### v1

# SIC-POVMs and orders of real quadratic fields

Abstract.
We consider the problem of counting and classifying symmetric informationally complete positive operator-valued measures (SICs or SIC-POVMs), that is, sets of $d^{2}$ equiangular lines in $\mathbb{C}^{d}$. For $4\leq d\leq 90$, we show the number of known equivalence classes of Weyl–Heisenberg covariant SICs in dimension $d$ equals the cardinality of the ideal class monoid of (not necessarily invertible) ideal classes in the real quadratic order of discriminant $(d+1)(d-3)$. Equivalently, this is the number of $\mathbf{GL}_{2}(\mathbb{Z})$ conjugacy classes in $\mathbf{SL}_{2}(\mathbb{Z})$ of trace $d-1$. We conjecture the equality extends to all $d\geq 4$. We prove that this conjecture implies more that one equivalence class of Weyl–Heisenberg covariant SICs for every $d>22$.
Additionally, we refine the “class field hypothesis” of Appleby, Flammia, McConnell, and Yard [10] to predict the exact class field generated by the ratios of vector entries for the equiangular lines defining a Weyl–Heisenberg covariant SIC. The class fields conjecturally associated to SICs in dimension $d$ come with a natural partial order under inclusion; we show that the natural inclusions of these fields are strict, except in a small family of cases.
Key words and phrases:
symmetric informationally complete positive operator-valued measure (SIC-POVM), complex equianglular lines, equiangular tight frame (ETF), real quadratic field, non-maximal order, binary quadratic form, ideal class group, explicit class field theory, Stark conjectures
1. Introduction
This paper concerns provable and conjectural properties of SIC-POVMs, which are ensembles of quantum measurements described by $d^{2}$ equiangular complex lines in $\mathbb{C}^{d}$. These properties are number-theoretic and relate to class numbers and class fields attached to orders of associated real quadratic number fields.
A set $S$ of complex lines through the origin in $\mathbb{C}^{d}$, written $S=\{\mathbb{C}\mathbf{v}_{1},\ldots,\mathbb{C}\mathbf{v}_{n}\}\subset\mathbb{P}^{d-1}(\mathbb{C})$, with $\mathbf{v}_{j}$ unit vectors, is equiangular if all the pairwise Hermitian inner products $\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle$ for $i\neq j$ have the same absolute value $|\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle|=\alpha\geq 0$. In 1975, Delsarte, Goethals, and Seidel [19] proved an upper bound of $n\leq d^{2}$ on the size of any set of complex equiangular lines in $\mathbb{C}^{d}$. They showed that if the bound $n=d^{2}$ is attained, then necessarily $\alpha=\frac{1}{d+1}$.
A symmetric informationally complete positive operator-valued measure or SIC-POVM (henceforth SIC) is a set of quantum measurements that is equivalent to a (maximal) set of $d^{2}$ complex equiangular lines in $\mathbb{C}^{d}$. Specifically, for a set of $d^{2}$ equiangular lines $\{\mathbb{C}\mathbf{v}_{1},\ldots,\mathbb{C}\mathbf{v}_{d^{2}}\}$ represented by unit column vectors $\mathbf{v}_{j}\in\mathbb{C}^{d}$, the associated SIC is the set of normalized rank one Hermitian projection operators ${\mathbf{\Pi}}[S]:=\left\{\frac{1}{d}\Pi_{1},\ldots,\frac{1}{d}\Pi_{d^{2}}\right\}$ where $\Pi_{j}=\mathbf{v}_{j}\mathbf{v}_{j}^{\dagger}$. (Vectors $\mathbf{v}$ are column vectors, $\mathbf{v}^{\dagger}$ is the conjugate transpose of $\mathbf{v}$, and the Hermitian inner product is $\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle=\mathbf{v}_{i}^{\dagger}\mathbf{v}_{j}$, following conventions from physics. Here, $\mathbf{v}_{i}^{\dagger}\mathbf{v}_{i}=1.$) The SIC condition is equivalent to the following $d^{4}$ identities on the traces of products of the projection operators:
$\operatorname{Tr}(\Pi_{i}\Pi_{j})$ =$=\begin{cases}1&\mbox{if }i=j,\\ \frac{1}{d+1}&\mbox{if }i\neq j,\end{cases}$ for $1\leq i,j\leq d^{2}$. (1.1)
We will use the term “SIC” interchangeably to mean either a “set of $d^{2}$ (scaled) Hermitian projection matrices ${\mathbf{\Pi}}(S)$ forming a SIC-POVM” or a “set of $d^{2}$ complex equiangular lines $S$ in $\mathbb{C}^{d}$.” When more precision is needed, we term the latter a line-SIC.
The name SIC-POVM was introduced in a 2004 paper of Renes, Blume-Kohout, Scott, and Caves [36], whose study was motivated by quantum state tomography, the problem of reconstructing a quantum state from a series of measurements performed on identical states. Scott showed that SICs are optimal measurement ensembles for tomography [37]; see also [26]. SICs have found further applications to quantum information processing [14, 42, 45, 18], quantum foundations [25], compressed sensing for radar [28], and classical phase retrieval [21].
The object of this paper is to extend the range of conjectures about the structure of SICs proposed by Appleby, Flammia, McConnell, and Yard [9]. Their conjectures relate Weyl–Heisenberg SICs in $\mathbb{C}^{d}$ (defined in Section 2) for $d\geq 4$ to class field theory over real quadratic fields. The ratios of entries of each of the $d^{2}$ vectors in such SICs lie in a number field that is an abelian extension of $K=\mathbb{Q}(\sqrt{\Delta_{d}})$ with $\Delta_{d}=(d+1)(d-3)$. Class field theory classifies the abelian extensions of a given number field $K$ in terms of ray class groups, which encode the arithmetic of ideals of the ring $\mathcal{O}_{K}$ of algebraic integers of $K$.
The number fields associated to individual SICs in dimension $d$ may vary with the SIC, and the subclass of SICs to which the most precise conjecture of [9] applies consists of those giving rise to minimal number field that occurs for a fixed $d$. Appleby, Flammia, McConnell, and Yard conjecture that this minimal field is a specific ray class field of $\mathbb{Q}(\sqrt{\Delta_{d}})$; see 1.1. The corresponding SICs comprise the minimal multiplet, as defined in Section 2. It is one of several multiplets, which are finite sets of geometric equivalence classes of SICs making up a single orbit of a Galois group action on geometric equivalence classes.
The current paper presents a refinement of these conjectures that accounts for all the remaining multiplets of Weyl–Heisenberg SICs, those having larger associated number fields. For $d\geq 4$ the refined conjecture says that each multiplet is indexed by a particular order $\mathcal{O}^{\prime}$ of a real quadratic number field $K_{d}=\mathbb{Q}(\sqrt{\Delta_{d}})$ intermediate between $\mathcal{O}=\mathbb{Z}[\varepsilon_{d}]$ with $\varepsilon_{d}=\frac{1}{2}(d-1+\sqrt{\Delta_{d}})$ and the maximal order $\mathcal{O}_{K_{d}}$. Here an order of a $K_{d}$ is a subring of $K_{d}$ whose additive group is free of rank $[K_{d}:\mathbb{Q}]=2$.
The number field associated to the $\mathcal{O}^{\prime}$-multiplet is conjectured to be a certain ray class field associated to the order $\mathcal{O}^{\prime}$. The definition of ray class fields associated to orders originates in the authors’ previous work [33]. Inclusions of orders induce reverse inclusion of the corresponding ray class fields, so the minimal multiplet corresponds to the maximal order.
1.1. SICs
Most work on SICs follows Zauner’s 1999 PhD thesis [48, 49] on quantum designs. A few sporadic examples (such as the Hoggar lines in dimension $8$ [29, 30]) were known prior to Zauner’s work. Zauner conjectured, based on his numerical investigations, that SICs exist in every dimension $d$, and moreover, that a SIC in $\mathbb{C}^{d}$ could be found as the orbit of a single vector under the Weyl–Heisenberg group [48, 49]; the latter objects are called Weyl–Heisenberg covariant SICs (or just Weyl–Heisenberg SICs) and are defined and discussed in Section 2. Weyl–Heisenberg SICs are the main topic of this paper. (An extensive treatment of the history of SICs is given by Fuchs, Hoang, and Stacy [24].)
Over the past 25 years, a large set of data has been computed about SICs in particular dimensions—sometimes approximately and sometime exactly—resulting in extensive datasets [39, 38, 22]. SICs have been rigorously constructed in a finite set of dimensions, including Weyl–Heisenberg SICs in all dimensions $1\leq d\leq 53$ and the Hoggar lines for $d=8$ (which are not Weyl–Heisenberg covariant). Beyond these exact constructions, numerical computations up to $d=90$ suggest there are only finitely many (equivalence classes of) Weyl–Heisenberg SICs in each dimension $d\neq 3$. It is known that $d=3$ has infinitely many equivalence classes of Weyl–Heisenberg SICs, and we will not consider it further.
The current numerical data is believed to account for all Weyl–Heisenberg SICs of dimension $4\leq d\leq 90$. There is a geometric action moving these SICs around by the normalizer of the Weyl–Heisenberg group inside the extended unitary group (which is a finite group whose order varies erratically with $d$, while the Weyl–Heisenberg group itself has order $d^{3}$.) In addition, since the (normalized) generating vectors of Weyl–Heisenberg SICs are (empirically) given by algebraic numbers, there is a Galois action, which sometimes moves SICs to new SICs and sometimes moves them to non-SICs.
The process of understanding this data has led to formulation of a series of interlocking hypotheses. The two most basic conjectures, which are both unsolved, are:
  1. (1)
    the existence of Weyl–Heisenberg SICs in all dimensions (conjectured by Zauner);
  2. (2)
    the finiteness of the set of (equivalence classes of) Weyl–Heisenberg SICs in every fixed dimension $d\neq 3$.
The finiteness assertion has been a “folklore” conjecture since about 2004 [2] and was observed informally up to dimension $d=50$ by Scott and Grassl (who state they are confident their list is complete for $d\leq 50$ in [39, Sec. 4]) and up to $d=90$ by Fuchs, Hoang, and Stacy [24, Sec. 6]. It was explicitly formulated as a conjecture in the first author’s PhD thesis [31, Conj. V.12]. A closely related conjecture was formulated by Waldron [46, Sec. 14.27].
The finiteness assertion would imply fiducial vectors of Weyl–Heisenberg SICs can be chosen to have all entries algebraic numbers. Further study of the algebraic numbers in existing examples has led to surprising discoveries.
1.2. Connection of SICs to algebraic number theory and class fields
Recent work has connected SICs to explicit class field theory over real quadratic number fields. Appleby, Yadsan-Appleby, and Zauner [4] discovered in 2013 that in all known examples, the algebraic number field $\mathbb{E}$ generated by the entries of the projections ${\mathbf{\Pi}}[S]$ contained the quadratic field $K_{d}=\mathbb{Q}(\sqrt{(d+1)(d-3)})$, and Galois automorphisms often mapped SICs to new SICs. Moreover, they observed that $\mathbb{E}$ was an abelian extension of $K_{d}$ in all known examples.
Such abelian Galois extensions are subfields of ray class fields over $K_{d}$. In 2016, Appleby, Flammia, McConnell, and Yard [10, 9] made explicit predictions relating the number fields generated by Weyl–Heisenberg SICs to ray class fields. Based on extensive data given in [10, Prop. 1], they formulated the following conjecture [10, Conj. 2].
Conjecture 1.1 (Appleby, Flammia, McConnell and Yard).
Let $d\geq 4$, $\Delta_{d}=(d+1)(d-3)$, and $K_{d}=\mathbb{Q}(\sqrt{\Delta_{d}})$.
  1. (1)
    At least one Weyl–Heisenberg line-SIC $S$ exists in dimension $d$ having SIC field $F^{\rm vec}\!\left(S\right)$ equal to $H_{d^{\prime}\infty_{1}\infty_{2}}$, the ray class field of $K_{d}$ having ray class modulus $d^{\prime}\infty_{1}\infty_{2}$, where $d^{\prime}=d$ if $d$ is odd, and $d^{\prime}=2d$ if $d$ is even.
  2. (2)
    Every Weyl–Heisenberg line-SIC $S$ in dimension $d$ has SIC-field $F^{\rm vec}\!\left(S\right)$ that is a finite Galois extension of $\mathbb{Q}$ containing $H_{d^{\prime}\infty_{1}\infty_{2}}$.
The technical definition of the SIC-field $F^{\rm vec}\!\left(S\right)$ appearing in 1.1 is given in Section 3.1. Ray class fields are discussed in Section 8.1.
Appleby, Flammia, McConnell, and Yard verified assertion (1) of 1.1 for
$d\in\{4,5,6,7,8,9,10,11,12,13,14,15,16,17,18,19,20,21,24,28,30,35,39,48\}.$ (1.2)
They verified assertion (2) of the conjecture for all currently known Weyl–Heisenberg SICs in these dimensions. They also verified that the fields $F^{\rm vec}\!\left(S\right)$ are all abelian over $K_{d}$ in the known cases. (The list of SICs used is believed to be complete in these dimensions.)
[…]
The origin of the discriminant $\Delta_{d}=(d+1)(d-3)$ appearing in the conjecture is not completely understood, but there are some tantalizing hints. The appearance of the factor $d+1$ dividing $\Delta_{d}$ seems related to the fact that in the inner products of distinct unit vectors in a SIC are of absolute value $\frac{1}{\sqrt{d+1}}$. The appearance of $d-3$ in the conjecture accounts for $d=3$ being special in that $\Delta_{3}=0$. The values $d=1,2$ have $\Delta_{d}<0$, so $K_{d}=\mathbb{Q}(\sqrt{\Delta_{d}})$ is an imaginary quadratic field, while for $d\geq 4$, $\Delta_{d}$ specifies a (nontrivial) real quadratic field. The occurrence of a small unit $\varepsilon_{d}=\frac{1}{2}(d-1+\sqrt{\Delta_{d}})$ (as compared to $\Delta_{d}$, which need not be a fundamental discriminant!) is certainly relevant. As noted by previously by the first author [32, p. 8], skipping repeated fields (e.g., $K_{8}=K_{4}$), the sequence $K_{4},K_{5},K_{6},K_{7},K_{9},K_{10},\ldots$ is precisely the ordering of all real quadratic fields by the size of the smallest nontrivial totally positive unit (which will necessarily be equal to $\varepsilon_{d}$).
1.3. Main results
In this paper, we refine 1.1 to conjectures that specify the number fields associated to all the Weyl–Heisenberg SICs in each dimension $d\geq 4$ and which count the number of such SICs associated to each field.
More precisely, we connect both SICs and fields to certain orders of the real quadratic number field $K_{d}=\mathbb{Q}(\sqrt{\Delta_{d}})$, with $\Delta_{d}=(d+1)(d-3)$. Write $\Delta_{d}=f^{2}\Delta_{0}$, where $\Delta_{0}$ is the fundamental discriminant $K_{d}$, and $f\geq 1$. The orders that arise are exactly those $\mathcal{O}^{\prime}$ satisfying $\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{K}$ for $\varepsilon_{d}=\frac{1}{2}(d-1+\sqrt{\Delta_{d}})$. These orders, arranged by set inclusion, form a distributive lattice, isomorphic to that of the positive divisors of $f$ under divisibility.
We organize Weyl–Heisenberg SICs into equivalence classes. For a line-SIC $S$,
  1. (1)
    the geometric equivalence class $[S]$ is the orbit of $S$ under the action of the projective extended Clifford group $\operatorname{PEC}(d)$ (see 2.9), and
  2. (2)
    the Galois multiplet $[[S]]$ is the set of geometric equivalence classes that are permuted with each other under a certain Galois action (see 4.11).
In Section 3, we define several fields attached to a line-SIC $S=\{\mathbb{C}\mathbf{v}_{k}:1\leq k\leq d^{2}\}$. These fields are finitely generated extensions of the rational numbers $\mathbb{Q}$.
  1. (1)
    The ratio SIC-field $F^{\rm vec}\!\left(S\right)$ (called the “SIC field” in [10]) is the field generated by all the ratios $\frac{v_{ki}}{v_{kj}}$ of distinct entries ($i\neq j$, $v_{kj}\neq 0$) of the vectors $\mathbf{v}_{k}$ with $\mathbf{v}_{k}\in S$.
  2. (2)
    The projection SIC-field $F^{\rm proj}\!\left(S\right)$ is the field generated by all the matrix entires of the Hermitian projection operators $\Pi_{i}=\mathbf{v}_{i}\mathbf{v}_{i}^{\dagger}$ of the $d^{2}$ unit vectors $\{{\mathbf{v}}_{i}:1\leq i\leq d^{2}\}$ defining $S$. It is shown to be the compositum of $F^{\rm vec}\!\left(S\right)$ with its complex conjugate field $\overline{F^{\rm vec}\!\left(S\right)}$ (in 3.3).
  3. (3)
    The extended projection SIC-field $F^{\rm eproj}\!\left(S\right)=F^{\rm proj}\!\left(S\right)\!(-e^{\frac{\pi i}{d}})$ is obtained by adjoining the root of unity $-e^{\frac{\pi i}{d}}$. For Weyl–Heisenberg SICs, we show $F^{\rm eproj}\!\left(S\right)$ is an invariant of the geometric equivalence class $[S]$ (in 3.8), and we show the index $[F^{\rm eproj}\!\left(S\right):F^{\rm proj}\!\left(S\right)]\in\{1,2\}$ and is $1$ if $d$ is odd (in 3.6).
  4. (4)
    For SICs defined over a finite extension of $\mathbb{Q}$, we define the Galois projection SIC-field $F^{\rm Gproj}\!\left(S\right)$ to be the normal closure of $F^{\rm eproj}\!\left(S\right)$, which is an invariant of the multiplet $[[S]]$. We show that the subgroup of $\operatorname{Gal}(F^{\rm Gproj}\!\left(S\right)/\mathbb{Q})$ consisting of elements that commute with complex conjugation sends SICs to other SICs with the same Galois projection SIC-field (in 4.8).
We have the field inclusions
$F^{\rm vec}\!\left(S\right)\subseteq F^{\rm proj}\!\left(S\right)\subseteq F^{\rm eproj}\!\left(S\right)\subseteq F^{\rm Gproj}\!\left(S\right).$ (1.3)
In 4.9, we show that 1.1 implies that
$F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)=F^{\rm eproj}\!\left(S\right)=F^{\rm Gproj}\!\left(S\right),$ (1.4)
whence $F^{\rm vec}\!\left(S\right)$ is a finite Galois extension of $\mathbb{Q}$.
The refinements of 1.1 formulated in this paper posit that Weil–Heisenberg SICs in dimension $d\neq 3$ occur in multiplets indexed by the intermediate orders $\mathcal{O}^{\prime}$ of $K_{d}$ the lie between $\mathcal{O}=\mathbb{Z}[\varepsilon_{d}]$ and $\mathcal{O}_{K_{d}}$. Each of these orders is in turn indexed by its conductor $f^{\prime}$, a positive integer such that $\operatorname{disc}(\mathcal{O}^{\prime})=(f^{\prime})^{2}\Delta_{0}$ where $\Delta_{0}=\operatorname{disc}(\mathcal{O}_{K_{d}})$. The conductors appearing are the positive divisors of the conductor $f$ of $\mathcal{O}$.
Conjecture 1.2 (Order-to-Multiplet Conjecture).
Fix a positive integer $d\neq 3$, let $\Delta=\Delta_{d}=(d+1)(d-3)$, and write $\Delta_{d}=f^{2}\Delta_{0}$ for $\Delta_{0}=\operatorname{disc}(\mathcal{O}_{K_{d}})$ and some positive integer $f$. Then there is a bijection $\mathcal{M}$ between conductors of intermediate orders and Galois multiplets,
[…]
having the following properties.
  • (1)
    The number of geometric equivalence classes $[S]$ in the multiplet $\mathcal{M}(f^{\prime})=[[S]]$ is $\left|\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\right|$, the class number of the invertible class group of the order $\mathcal{O}^{\prime}$ of discriminant $(f^{\prime})^{2}\Delta_{0}$.
  • (2)
    If $f_{1}\mid f_{2}\mid f$, $[S_{1}]\in\mathcal{M}(f_{1})$, and $[S_{2}]\in\mathcal{M}(f_{2})$, then $F^{\rm vec}\!\left(S_{1}\right)\subseteq F^{\rm vec}\!\left(S_{2}\right)$.
A mildly strengthened version of this conjecture is stated as 7.2. The field inclusion property (2) puts a strong restriction on the map $\mathcal{M}$. Modulo the next conjecture, it determines $\mathcal{M}$ entirely for a fraction $\frac{47}{48}$ of all dimensions $d$ (in the sense of 9.10).
The following conjecture is an extension of 1.1 using the notion of ray class fields of an order $\mathcal{O}^{\prime}$ introduced in [33] and reviewed in Section 8.1.
Conjecture 1.3 (Ray Class Fields of Orders Conjecture).
Let $d\geq 4$. Then for any Weyl–Heisenberg line-SIC $S$ in dimension $d$ with associated order $\mathcal{O}^{\prime}$ having discriminant $(f^{\prime})^{2}\Delta_{0}$ (that is, with $\mathcal{M}(f^{\prime})=[[S]]$),
$F^{\rm vec}\!\left(S\right)=H_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}$ (1.6)
where $H_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}$ is the ray class field having level datum $(\mathcal{O}^{\prime};d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\})$, $d^{\prime}=2d$ if $d$ is even, and $d^{\prime}=d$ if $d$ is odd.
The special case $f^{\prime}=1$ corresponds to 1.1, where the order $\mathcal{O}^{\prime}$ is the maximal order. The field inclusions asserted in 1.1(2) follow from 1.2. Appleby, Flammia, McConnell, and Yard [9] call the associated Galois multiplet $[[S]]$ the minimal multiplet, as it has the smallest associated field $F^{\rm vec}\!\left(S\right)$ under inclusion. 8.4 gives a further refinement of 1.3; these conjectures refine and revise [31, Conj. V.17].
The paper describes numerical checks of predictions of these conjectures against numerical data. The data for $4\leq d\leq 90$ exhibits patterns consistent with grouping SICs into multiplets indexed by orders. We formulate several narrower conjectures that make numerical predictions that can be checked against existing data and give tables of comparisons.
We also derive some consequences of the refined conjectures. In Section 6, we show they imply that the number of distinct geometric equivalence classes $[S]$ will be larger than $1$ for all $d\geq 23$ and will grow like $d^{1+o(1)}$ as $d\to\infty$. (See 6.3 and 6.4.)
1.4. Notation
  • Boldface roman upper case letters represent continuous groups, e.g., $\mathbf{EU}(d)$.
  • Boldface roman lower case letters represent column vectors in $\mathbb{C}^{d}$. For a complex vector, $\mathbf{w}^{\dagger}:=\overline{\mathbf{w}}^{T}$, and similarly for matrices.
  • Upper case roman letters represent $d\times d$ complex matrices (e.g., $X$, $Z$); they also appear in names of finite matrix groups (e.g., $\operatorname{WH}(d)$) and sets (e.g., $\operatorname{WHSIC}(d)$).
  • Matrices and vectors have rows and columns numbered from 00 to $d-1$. We follow the text of Waldron [46], which identifies indices with $\mathbb{Z}/d\mathbb{Z}$.
  • Greek letters generally represent algebraic numbers (e.g., $\xi_{d}=-e^{\frac{\pi i}{d}}$ and $\zeta_{d}=e^{\frac{2\pi i}{d}}$).
  • However, the letter $\sigma$ is exclusively used for Galois automorphisms, with $\sigma_{c}$ denoting complex conjugation. Galois group actions are treated either as acting on scalars or as acting entrywise on vectors or matrices.
  • The notation $S$ is used for a ($d$-dimensional) line-SIC; it is a set of $d^{2}$ complex lines.
  • The notation $[S]$ is used for a geometric equivalence class of SICs. (However, $[\mathbf{v}]$ denotes the complex line $\mathbb{C}\mathbf{v}$, when $\mathbf{v}$ is a vector.)
  • We use the convention that $d^{\prime}$ means $d^{\prime}=d$ if $d$ is odd and $d^{\prime}=2d$ if $d$ is even. The number $\xi_{d}=-e^{\frac{\pi i}{d}}$ is a $d^{\prime}$-th root of unity.
  • We write $\Delta_{d}=(d+1)(d-3)$ for the (often non-fundamental) discriminant associated to dimension $d$, $K_{d}=\mathbb{Q}(\sqrt{\Delta_{d}})$ for the associated field, and $\varepsilon_{d}=\frac{1}{2}(d-1+\sqrt{\Delta_{d}})$ for the associated unit.
2. Weyl–Heisenberg SICs
In this section, we defined and give background on Weyl–Heisenberg (covariant) SICs, which satisfy a particular group symmetry property.
2.1. Weyl–Heisenberg SICs
Specifically, Weyl–Heisenberg line-SICs will be defined as the orbit of a single line under the action of the Weyl-Heisenberg group, a finite subgroup of the group of $d\times d$ unitary matrices.
Definition 2.1.
The Weyl–Heisenberg group is $\operatorname{WH}(d)=\{\xi^{r}X^{p}Z^{q}:p,q,r\in\mathbb{Z}\}$, where $\xi=-e^{\frac{\pi i}{d}}$, $\zeta=\xi^{2}=e^{\frac{2\pi i}{d}}$, and
$X$ $=\left(\begin{array}[]{ccccc}0&0&\cdots&0&1\\ 1&0&\cdots&0&0\\ 0&1&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&1&0\end{array}\right),$ $Z$ $=\left(\begin{array}[]{ccccc}1&0&0&\cdots&0\\ 0&\zeta&0&\cdots&0\\ 0&0&\zeta^{2}&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&\zeta^{d-1}\end{array}\right),$ (2.11)
The projective Weyl–Heisenberg group $\operatorname{PWH}(d)$ is defined to be $\operatorname{WH}(d)$ modulo its scalar matrix subgroup $Z(\operatorname{WH}(d))=\{\xi^{j}I:\,j\geq 0\}$, where $I$ is the $d\times d$ identity matrix.
It is easily checked that $ZX=\zeta XZ$. It follows that $\operatorname{WH}(d)$ is a group. Its order is $d^{3}$ if $d$ is odd and $2d^{3}$ if $d$ is even, noting that $\xi$ is a $d$-th root of unity for odd $d$ and a $2d$-th root of unity for even $d$. The center $\operatorname{Z}(\operatorname{WH}(d))=\langle\xi I\rangle$ of $\operatorname{WH}(d)$ is given by scalar matrices, and the quotient group $\operatorname{PWH}(d)\cong\operatorname{WH}(d)/\operatorname{Z}(\operatorname{WH}(d))\cong\left(\mathbb{Z}/d\mathbb{Z}\right)^{2}$.
Definition 2.2.
For $\mathbf{w}\in\mathbb{C}^{d}$, its Weyl–Heisenberg orbit is
$\operatorname{orb}_{\operatorname{WH}}(\mathbf{w}):=\{H\mathbf{w}:\,H\in\operatorname{WH}(d)\}.$ (2.12)
Its associated set of complex lines, denoted $\operatorname{WHL}(\mathbf{w})$ is
$\operatorname{WHL}(\mathbf{w}):=\{[\mathbf{w}]:\,\mathbf{w}\in\operatorname{orb}_{\operatorname{WH}}(\mathbf{w})\}=\{[H\mathbf{w}]:H\in\operatorname{PWH}(d)\}.$ (2.13)
For any $\mathbf{w}$, $\operatorname{WHL}(\mathbf{w})$ has cardinality at most $d^{2}$, and for generic $\mathbf{w}$, exactly $d^{2}$.
Definition 2.3.
A Weyl–Heisenberg covariant SIC (or simply a Weyl–Heisenberg SIC) is any $S(\mathbf{v}):=\operatorname{WHL}(\mathbf{v})$ for which ${\mathbf{\Pi}}(S)$ is a $d$-dimensional SIC. We denote the set of all Weyl–Heisenberg SICs in dimension $d$ as $\operatorname{WHSIC}(d)$.
We call a generating vector $\mathbf{v}$ of a Weyl–Heisenberg SIC a fiducial vector. Such a vector is well-defined up to multiplication by a complex scalar $c\in\mathbb{C}^{\times}$. Each SIC has a choice of $d^{2}$ fiducial vectors, up to scalars. We call two Weyl–Heisenberg SICs scalar-equivalent if they have a common fiducial vector, up to multiplication by a nonzero scalar.
Conjecture 2.4 (Strong Zauner Conjecture).
For each dimension $d\geq 1$, there exists at least one Weyl–Heisenberg SIC.
This conjecture is currently known, by explicit constructions to hold for finitely many dimensions $d$. At present it includes dimensions $1\leq d\leq 53$ and a extensive number of higher dimensions.
Conjecture 2.5 (Weyl–Heisenberg Boundedness Conjecture).
For each dimension $d\geq 1$, except $d=3$, there are finitely many scalar-equivalence classes of Weyl–Heisenberg SICs.
There are a few small dimensions where this conjecture has been proved rigorously, including $d\in\{1,2,4,5\}$. In higher dimensions, up to at least $d=90$, it is supported by non-rigorous evidence given by numerical optimization routines repeatedly returning the same SICs [39, 38]. In dimension $3$, there is a one-parameter family of inequivalent Weyl–Heisenberg SICs giving uncountably many distinct scalar-equivalence classes of SICs.
A closely related conjecture is stated in Waldron [46, Sec. 14.27]: The (real) algebraic variety of (normalized Weyl–Heisenberg) SIC fiducials is zero-dimensional, except for $d=3$, where it is one-dimensional. This conjecture implies 2.5; it is equivalent if one takes “dimension” to mean “real dimension” (i.e., dimension of the Zariski closure) and is a stronger conjecture if one takes “dimension” to mean “dimension of the complex variety defined by the same equations.” Resolving 2.5 or Waldron’s conjecture in general (and specifically even for $n=6$) is an interesting open problem in optimization and real algebraic geometry.
2.2. Geometric equivalence classes of Weyl–Heisenberg SICs
There are two types of geometric operators acting on $\mathbb{C}^{d}$ that send SICs to SICs. Firstly, if $U$ is a unitary matrix in $\mathbf{U}(d)=\{U\in\mathbf{GL}(\mathbb{C}^{d}):UU^{\dagger}=1\}$ and $S=\{\mathbb{C}\mathbf{v}_{1},\ldots,\mathbb{C}\mathbf{v}_{d^{2}}\}$ is a SIC, then so is $US=\{\mathbb{C}U\mathbf{v}_{1},\ldots,\mathbb{C}U\mathbf{v}_{d^{2}}\}$. (This is a unitary equivalence.) Secondly, if $C_{d}$ is the complex conjugation operator acting pointwise on $\mathbb{C}^{d}$, so that $C_{d}\mathbf{v}:=\overline{\mathbf{v}}$, then $C_{d}$ also preserves the SIC property, as does any “antiunitary” operator of the form $C_{d}U$. These operators may be collected together to form the extended unitary group.
Definition 2.6.
The extended unitary group is $\mathbf{EU}(d):=\mathbf{U}(d)\sqcup C_{d}\mathbf{U}(d)$. The projective extended unitary group $\mathbf{PEU}(d)$ is defined to be $\mathbf{EU}(d)$ modulo its scalar matrix subgroup $\{zI:z\in\mathbb{C},\left|z\right|=1\}$.
All known SICs are group covariant, meaning that they are the orbit of a vector under the action of a finite subgroup of the unitary group.
Definition 2.7.
A SIC $S$ is group covariant for a subgroup $G$ of $\mathbf{U}(d)$ if there exists a vector $v\in\mathbb{C}^{d}$, called a fiducial vector, such that $S=\{\mathbb{C}Av:A\in G\}$.
If $\mathbf{v}$ is a fiducial vector for a Weyl–Heisenberg covariant SIC and $U\in\mathbf{EU}(d)$, then the vector $U\mathbf{v}$ will be a fiducial vector for a SIC covariant for the group $G^{\prime}=U\operatorname{WH}(d)U^{-1}$. All currently known SICs are unitary-equivalent to a Weyl–Heisenberg group $\operatorname{WH}(d)$ covariant SIC, with a single exception, the Hoggar lines in $\mathbb{C}^{8}$ [29, 30]. The Hoggar lines are group covariant for the tensor product $\operatorname{WH}(2)\otimes\operatorname{WH}(2)\otimes\operatorname{WH}(2)$. (The group $\operatorname{WH}(2)\otimes\operatorname{WH}(2)\otimes\operatorname{WH}(2)$ may itself be thought of as a generalized (Weyl)-Heisenberg group.)
The extended unitaries $U\in\mathbf{EU}(d)$ that preserve the Weyl–Heisenberg covariance property are precisely those for which $U\operatorname{WH}(d)U^{-1}=\operatorname{WH}(d)$, that is, elements of the extended Clifford group, defined as follows.
Definition 2.8.
We define several closely related groups.
  • (1)
    The Clifford group $\operatorname{C}(d)$ is the normalizer of $\operatorname{WH}(d)$ in $\mathbf{U}(d)$.
  • (2)
    The projective Clifford group $\operatorname{PC}(d)$ is the quotient group of $\operatorname{C}(d)$ modulo its scalar matrix subgroup.
  • (3)
    The extended Clifford group $\operatorname{EC}(d)$ is the normalizer of $\operatorname{WH}(d)$ in $\mathbf{EU}(d)$.
  • (4)
    The projective extended Clifford group $\operatorname{PEC}(d)$ is the quotient of the extended Clifford group $\operatorname{EC}(d)$ modulo its scalar matrix subgroup.
The Clifford group $\operatorname{C}(d)$ is infinite because it contains a copy of $\mathbf{U}(1)$ as scalar matrices $zI$ with $\left|z\right|=1$; thus, $\operatorname{EC}(d)$ is infinite. However, $\operatorname{PC}(d)$ and $\operatorname{PEC}(d)$ are finite groups. (The extended Clifford $\operatorname{EC}(d)$ group is studied in connection with SICs by Appleby [1].)
Definition 2.9.
The set of geometric equivalence classes of Weyl–Heisenberg SICs is the set of equivalence classes of $\operatorname{WHSIC}(d)$ under the equivalence relation $[S({\mathbf{v}})]\sim[S(U\mathbf{v})]$ whenever $U\in\operatorname{PEC}(d)$ (equivalently, whenever $U\in\operatorname{EC}(d)$.) This set of equivalence classes is denoted $\operatorname{WHSIC}(d)/\operatorname{PEC}(d)$. (It is not a group.) We denote the geometric equivalence class of a Weyl–Heisenberg SIC $S$ by $[S]$.
3. Fields generated by SICs
In this section, we define several number fields attached to arbitrary SICs. Many of these fields have been studied before, but they are often discussed within the context of assuming various conjectures about SICs. Here, we focus on relationships between “SIC fields” that can be proven rigorously, either in general (Section 3.1) or in the case of Weyl–Heisenberg SICs (Section 3.2).
In Section 4, we will focus on SICs expressible in terms of algebraic numbers and discuss some conditional results about fields generated by SICs in that setting.
3.1. Fields generated by general SICs
The fields we will study here are each associated to an individual SIC $S$ and depend only on data computed from the orthogonal projections $\{\Pi_{i}:1\leq i\leq d^{2}\}$ defining the SIC.
Definition 3.1.
Let ${\mathbf{\Pi}}(S)=\{\frac{1}{d}\Pi_{1},\ldots,\frac{1}{d}\Pi_{d^{2}}\}$ be an arbitrary SIC in dimension $d$, with each $\Pi_{i}=\mathbf{v}_{i}\mathbf{v}_{i}^{\dagger}$ for a unit vector $\mathbf{v}_{i}$. (Thus each $\mathbf{v}_{i}$ is a unit fiducial vector, unique up to multiplication by a unit scalar $e^{i\theta}$.)
  • (1)
    The ratio SIC-field $F^{\rm vec}\!\left(S\right)$ is the field extension of $\mathbb{Q}$ generated by all the ratios $v_{ij}/v_{ik}$ (for $v_{ik}\neq 0$) of the entries of the $d^{2}$ fiducial vectors $\mathbf{v}_{i}$ of $S$.
  • (2)
    The triple product SIC-field $F^{\rm trip}\!\left(S\right)$ is the field extension of $\mathbb{Q}$ generated by all the “triple products” $\operatorname{Tr}(\Pi_{i}\Pi_{j}\Pi_{k})$.
  • (3)
    The unitary invariant SIC-field $F^{\rm inv}\!\left(S\right)$ is the field extension of $\mathbb{Q}$ generated by all the unitary invariants $\operatorname{Tr}(\Pi_{i_{1}}\Pi_{i_{2}}\cdots\Pi_{i_{n}})$ for $n\in\mathbb{N}$.
  • (4)
    The projection SIC-field $F^{\rm proj}\!\left(S\right)$ is the field extension of $\mathbb{Q}$ generated by all the entries of the $d^{2}$ Hermitian projection matrices $\Pi_{j}$.
  • (5)
    The extended projection SIC-field $F^{\rm eproj}\!\left(S\right):=F^{\rm proj}\!\left(S\right)\!(\xi_{d})$, adjoining $\xi_{d}=-e^{\frac{\pi i}{d}}$.
A priori, these fields could be transcendental extensions of $\mathbb{Q}$, and this does occur in dimension $d=3$.
Remark 3.2.
These fields have previously been studied with various names and notations, which we now review.
  1. (1)
    The field $F^{\rm vec}\!\left(S\right)$ is the SIC field $\mathbb{Q}[S]$ attached to a general SIC $S$ by Appleby, Flammia, McConnell, and Yard [10, p. 212 bottom], as used in Thm. 1 and Conj. 2 of that paper. By definition, this field is independent of scaling factors in the $\mathbf{v}_{i}$. It is determined by ${\mathbf{\Pi}}(S)$ because the ratios
    $\frac{v_{ij}}{v_{ik}}=\frac{v_{ij}\overline{v}_{ik}}{v_{ik}\overline{v}_{ik}}=\frac{(\Pi_{i})_{jk}}{(\Pi_{i})_{kk}}.$ (3.1)
    We may rescale each of the $d^{2}$ vectors in $S$ to give $S=\{\mathbb{C}{\mathbf{w}}_{i}:1\leq i\leq d^{2}\}$ so that all entries of each ${\mathbf{w}}_{i}$ belong to $F^{\rm vec}\!\left(S\right)$. A given ${\mathbf{v}}_{i}=\left(v_{i0},v_{i1},\cdots,v_{i(d-1)}\right)^{T}$ has at least one nonzero entry, say $v_{ik}$, and we may set ${\mathbf{w}}_{i}=\frac{1}{v_{ik}}{\mathbf{v}}_{i}$, whence $w_{ij}=v_{ij}/v_{ik}\in F^{\rm vec}\!\left(S\right)$. The vectors ${\mathbf{w}}_{i}$ obtained this way may not be unit vectors.
    Prop. 3 of [10] states that, for a Weyl–Heisenberg SIC, the ratio SIC-field is the same for any SIC $S^{\prime}\in[S]$. The presented proof outline [10, Sec. 7.1] appears to assume some conjectural equivalences and characterizations of other related fields.
  2. (2)
    Concerning the triple product SIC-field $F^{\rm trip}\!\left(S\right)$, the fact that a SIC is determined up to unitary equivalence by its triple products was proved in 2011 by Appleby, Flammia, and Fuchs [3, Thm. 3]. We show unconditionally that $F^{\rm trip}\!\left(S\right)$ is an invariant of the geometric $\operatorname{PEC}(d)$-orbit of a Weyl–Heisenberg SIC $S$; see 3.4.
  3. (3)
    The unitary invariant SIC-field $F^{\rm inv}\!\left(S\right)$ is generated by the full set of projective unitary invariants of a general finite set of lines in $\mathbb{C}^{d}$; see [15] and [46, Ch. 8, Thm. 8.2]. We show in 3.3 that $F^{\rm trip}\!\left(S\right)=F^{\rm inv}\!\left(S\right)$.
  4. (4)
    The projection SIC-field $F^{\rm proj}\!\left(S\right)$, in the special case that $S$ is a Weyl–Heisenberg SIC, coincides with the field $L$ appearing in [10, Sec. 7], defined as the smallest subfield of $\mathbb{C}$ generated by the entries $v_{j}\overline{v}_{k}$ of unit vectors generating $\Pi_{\mathbf{v}}$ taken over all $d^{2}$ projectors $\Pi_{\mathbf{v}}$ of a Weyl–Heisenberg SIC.
  5. (5)
    The extended projection SIC-field $F^{\rm eproj}\!\left(S\right)$ is the SIC field $\mathbb{E}$ as defined by Appleby, Flammia, McConnell, and Yard [9, Sec. 4] for any SIC projector $S$. This definition of the SIC field is also used in [7]. Appleby, Flammia, McConnell, and Yard [9] state (without proof) that the field $\mathbb{E}$ is well-defined on the $\operatorname{PEC}(d)$-orbit $[S]$ of a Weyl–Heisenberg SIC $S$. We prove so unconditionally in 3.8.
The next result establishes some inclusion relations among these fields.
Proposition 3.3.
Let $S$ be an arbitrary line-SIC in dimension $d$. The fields associated to $S$ are related in the following ways.
  • (1)
    $F^{\rm proj}\!\left(S\right)=F^{\rm vec}\!\left(S\right)\overline{F^{\rm vec}\!\left(S\right)}$ (the compositum of $F^{\rm vec}\!\left(S\right)$ and its complex conjugate field). In particular, $F^{\rm proj}\!\left(S\right)$ is invariant under complex conjugation.
  • (2)
    $F^{\rm trip}\!\left(S\right)=F^{\rm inv}\!\left(S\right)\subseteq F^{\rm proj}\!\left(S\right)$.
Proof.
Write ${\mathbf{\Pi}}(S)=\{\frac{1}{d}\Pi_{1},\ldots,\frac{1}{d}\Pi_{d^{2}}\}$.
Proof of (1): Write $\Pi_{i}=\mathbf{v}_{i}\mathbf{v}_{i}^{\dagger}$, where $\mathbf{v}_{i}$ has entries $v_{ij}$ and $\left|\mathbf{v}_{i}\right|^{2}=1$. The nonzero entries of $\Pi_{i}$ are of the form $v_{ij}\overline{v}_{ik}$ for $v_{ij}\neq 0$, $v_{ik}\neq 0$ These entries can be expressed as
$v_{ij}\overline{v}_{ik}=\frac{v_{ij}\overline{v}_{ik}}{\left|\mathbf{v}_{i}\right|^{2}}=\frac{1}{\sum_{\ell=1}^{d}(v_{i\ell}/v_{ij})\overline{(v_{i\ell}/v_{ik})}}\in F^{\rm vec}\!\left(S\right)\overline{F^{\rm vec}\!\left(S\right)}.$ (3.2)
Thus, $F^{\rm proj}\!\left(S\right)\subseteq F^{\rm vec}\!\left(S\right)\overline{F^{\rm vec}\!\left(S\right)}$. To show the reverse inclusion, note that for $v_{ik}\neq 0$,
$v_{ij}/v_{ik}=\frac{v_{ij}\overline{v}_{ik}}{v_{ik}\overline{v}_{ik}}\mbox{ and }\overline{(v_{ij}/v_{ik})}=\frac{v_{ik}\overline{v}_{ij}}{v_{ik}\overline{v}_{ik}}.$ (3.3)
Proof of (2): First we show the equality $F^{\rm trip}\!\left(S\right)=F^{\rm inv}\!\left(S\right)$. Clearly $F^{\rm trip}\!\left(S\right)\subseteq F^{\rm inv}\!\left(S\right)$. To show the reverse inclusion, note that the $\Pi_{i}$ form a basis for the space of $d\times d$ complex matrices, and express
$\Pi_{i}\Pi_{j}=\sum_{\ell}\alpha_{ij}^{\ell}\Pi_{\ell}$ (3.4)
for structure constants $\alpha_{ij}^{\ell}\in\mathbb{C}$. We have
$\sum_{\ell}\alpha_{ij}^{\ell}=\operatorname{Tr}(\Pi_{i}\Pi_{j})=\frac{\delta_{ij}d+1}{d+1},\mbox{ where }\delta_{ij}=\begin{cases}1,&\mbox{if }i=j,\\ 0,&\mbox{if }i\neq j.\end{cases}$ (3.5)
Moreover, the triple products are related to the structure constants by
$\operatorname{Tr}(\Pi_{i}\Pi_{j}\Pi_{k})$ $=\sum_{\ell}\alpha_{ij}^{\ell}\operatorname{Tr}(\Pi_{k}\Pi_{\ell})$ (3.6)
$=\frac{d}{d+1}\alpha_{ij}^{k}+\frac{1}{d+1}\sum_{\ell}\alpha_{ij}^{\ell}$ (3.7)
$=\frac{d}{d+1}\alpha_{ij}^{k}+\frac{\delta_{ij}d+1}{(d+1)^{2}}.$ (3.8)
Thus, $\alpha_{ij}^{k}=\frac{d+1}{d}\operatorname{Tr}(\Pi_{i}\Pi_{j}\Pi_{k})-\frac{d(\delta_{ij}d+1)}{d+1}$, and the field generated over $\mathbb{Q}$ by the structure constants is $F^{\rm trip}\!\left(S\right)$. Moreover, using (3.4) repeatedly, we can express any product $\Pi_{i_{1}}\Pi_{i_{2}}\cdots\Pi_{i_{n}}$ in the $\left\{\Pi_{\ell}\right\}_{\ell}$-basis with coefficients that are integer polynomials in the structure constants. Thus, since $\operatorname{Tr}(\Pi_{\ell})=1$, $\operatorname{Tr}(\Pi_{i_{1}}\Pi_{i_{2}}\cdots\Pi_{i_{n}})$ is also an integer polynomials in the structure constants and is thus contained in $F^{\rm trip}\!\left(S\right)$.
Finally, the inclusion $F^{\rm inv}\!\left(S\right)\subseteq F^{\rm proj}\!\left(S\right)$ of (2) follows because $\operatorname{Tr}(\Pi_{i_{1}}\Pi_{i_{2}}\cdots\Pi_{i_{n}})$ is an integer polynomial in the entries of the various $\Pi_{j}$. ∎
Proposition 3.4.
Let $S$ be an arbitrary line-SIC in dimension $d$. Then the triple product field satisfies the invariance property $F^{\rm trip}\!\left(US\right)=F^{\rm trip}\!\left(S\right)$ for any $U\in\mathbf{EU}(d)$.
Proof.
Let $S=\{\mathbb{C}{\mathbf{v}}_{i}:1\leq i\leq d^{2}\}$ be an arbitrary line-SIC, and set ${\mathbf{\Pi}}(S)=\{\frac{1}{d}\Pi_{1},\ldots,\frac{1}{d}\Pi_{d^{2}}\}$, with $\Pi_{i}={\mathbf{v}}_{i}{\mathbf{v}}_{i}^{\dagger}$. Then ${\mathbf{\Pi}}(US)=\{\frac{1}{d}\Pi_{1}^{\prime},\ldots,\frac{1}{d}\Pi_{d^{2}}^{\prime}\}$. Write $U=CV$, where $C$ is either the identity or complex conjugation, and $V\in\mathbf{U}(d)$; then either $\Pi_{i}^{\prime}=V^{-1}\Pi_{i}V$, or $\Pi_{i}^{\prime}=V^{-1}\Pi_{i}^{\top}V$. In the first case, $\operatorname{Tr}(\Pi_{i}^{\prime}\Pi_{j}^{\prime}\Pi_{k}^{\prime})=\operatorname{Tr}(V^{-1}\Pi_{i}\Pi_{j}\Pi_{k}V)=\operatorname{Tr}(\Pi_{i}\Pi_{j}\Pi_{k})$. In the second case, $\operatorname{Tr}(\Pi_{i}^{\prime}\Pi_{j}^{\prime}\Pi_{k}^{\prime})=\operatorname{Tr}\!\left(V^{-1}\Pi_{i}^{\top}\Pi_{j}^{\top}\Pi_{k}^{\top}V\right)=\operatorname{Tr}\!\left((\Pi_{k}\Pi_{j}\Pi_{i})^{\top}\right)=\operatorname{Tr}(\Pi_{k}\Pi_{j}\Pi_{i})$. So $F^{\rm trip}\!\left(US\right)=F^{\rm trip}\!\left(S\right)$ in both cases. ∎
We derive some unconditional implications when these fields give Galois extensions of $\mathbb{Q}$.
Lemma 3.5.
Let $S$ be an arbitrary line-SIC in dimension $d$.
  • (1)
    If $F^{\rm vec}\!\left(S\right)$ is invariant under complex conjugation $\sigma_{c}$, then $F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right).$ Thus, if $F^{\rm vec}\!\left(S\right)$ is a Galois extension of $\mathbb{Q}$, then $F^{\rm proj}\!\left(S\right)$ is a Galois extension of $\mathbb{Q}$.
  • (2)
    If $F^{\rm proj}\!\left(S\right)$ is a Galois extension of $\mathbb{Q}$, then $F^{\rm eproj}\!\left(S\right)$ is a Galois extension of $\mathbb{Q}$.
Proof.
Proof of (1): The implication $F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)$ follows from 3.3(1).
Proof of (2): If $K=F^{\rm proj}\!\left(S\right)$ is Galois over $\mathbb{Q}$, then $F^{\rm eproj}\!\left(S\right)$ is the compositum $KL$, with $L=\mathbb{Q}(e^{\frac{\pi i}{d}})$ an abelian extension of $\mathbb{Q}$, so the compositum is a Galois extension of $\mathbb{Q}$. ∎
3.2. Fields generated by Weyl–Heisenberg SICs
We now derive some properties of the field $F^{\rm vec}\!\left(S\right)$, $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$ in the special case of Weyl–Heisenberg SICs.
Lemma 3.6.
Let $S$ be a Weyl–Heisenberg line-SIC $S=S([{\mathbf{v}}])$ of dimension $d$. Then:
  • (1)
    $F^{\rm proj}\!\left(S\right)$ contains a $d$-th root of unity $\zeta_{d}=e^{\frac{2\pi i}{d}}$.
  • (2)
    The index $[F^{\rm eproj}\!\left(S\right):F^{\rm proj}\!\left(S\right)]\in\{1,2\}$. It is $1$ whenever $d$ is odd.
Proof.
Proof of (1): Let ${\mathbf{v}}=(v_{0},v_{1},\ldots,v_{d-1})^{\top}$ be a fiducial vector for $S$, and let $\Pi={\mathbf{v}}{\mathbf{v}}^{\dagger}$. It has $(1,0)$-entry $(\Pi)_{10}=v_{1}\overline{v}_{0}\in F^{\rm proj}\!\left(S\right)$. Now the modulation matrix $Z\in\operatorname{WH}(d)$ gives another SIC vector $Z{\mathbf{v}}=(v_{0},e^{\frac{2\pi i}{d}}v_{1},\cdots,e^{\frac{2\pi i(d-1)}{d}}v_{d-1})$. The corresponding projection $\Pi^{\prime}:=Z{\mathbf{v}}(Z{\mathbf{v}})^{\dagger}=Z{\mathbf{v}}{\mathbf{v}}^{\dagger}Z^{-1}$ has $(\Pi^{\prime})_{1,0}=e^{\frac{2\pi i}{d}}v_{1}\overline{v}_{0}\in F^{\rm proj}\!\left(S\right)$. If $v_{1}\overline{v}_{0}\neq 0$, then we may conclude $e^{\frac{2\pi i}{d}}\in F^{\rm proj}\!\left(S\right)$.
We show every Weyl–Heisenberg SIC has a fiducial vector ${\mathbf{v}}$ with both $v_{0}$ and $v_{1}$ nonzero, so the argument applies. We first claim that every fiducial vector ${\mathbf{v}}$ has two consecutive entries $v_{i},v_{i+1}$ nonzero, allowing the wraparound product $v_{d-1}v_{0}$. Suppose not; then, the translation matrix $X\in\operatorname{WH}(d)$ gives a cyclic shift $X{\mathbf{v}}=(v_{1},v_{2},...,v_{d-1},v_{0})$, so $\langle{\mathbf{v}},X{\mathbf{v}}\rangle=\sum_{i=0}^{d-1}v_{i}\overline{v}_{i+1}=0$, contradicting the SIC property $\left|\langle{\mathbf{v}},X{\mathbf{v}}\rangle\right|=\frac{1}{d+1}$. The claim follows. If the given ${\mathbf{v}}$ has $v_{i}v_{i+1}\neq 0$, then ${\mathbf{v}}^{\prime}=X^{d-i}{\mathbf{v}}$ cyclically shifts entries so that $v_{0}^{\prime}v_{1}^{\prime}\neq 0$.
Proof of (2): By (1), $F^{\rm eproj}\!\left(S\right)=F^{\rm proj}\!\left(S\right)(e^{\frac{\pi i}{d}})$ is at most a quadratic extension of $F^{\rm proj}\!\left(S\right)$, since the field index $[\mathbb{Q}(e^{\frac{\pi i}{d}}):\mathbb{Q}(e^{\frac{2\pi i}{d}})]=2$ if $d$ is even and is $1$ if $d$ is odd. ∎
We now study the field $F^{\rm eproj}\!\left(S\right)$ on the $\operatorname{PEC}(d)$-orbit of $S$ and show it is a $\operatorname{PEC}(d)$-invariant. We recall a characterization of generators of the Clifford group.
Proposition 3.7.
The Clifford group $\operatorname{C}(d)$ is generated by all unitary scalar matrices $zI$ with $\left|z\right|=1$, together with
$X,Z,F,R$ (3.9)
where $X,Z$ are Weyl–Heisenberg group generators, $F$ is the discrete Fourier transform,
$F_{j,k}=\frac{1}{\sqrt{d}}e^{\frac{2\pi ijk}{d}}\,\mbox{ for }\,0\leq j,k\leq d-1.$ (3.10)
and $R$ is a diagonal matrix, with diagonal entries
$R_{j,j}=\xi_{d}^{j(j+d)}=(-1)^{jd}e^{\frac{\pi ij^{2}}{d}}\,\mbox{ for }\,0\leq j\leq d-1.$ (3.11)
Proof.
This is [46, Thm. 14.1]. ∎
Proposition 3.8.
Let $S$ be a Weyl–Heisenberg line-SIC $S=S([{\mathbf{v}}])$ of dimension $d$. For any $M\in\operatorname{EC}(d)$,
$F^{\rm eproj}\!\left(MS\right)=F^{\rm eproj}\!\left(S\right),$ (3.12)
where $MS$ denotes the Weyl–Heisenberg line-SIC having fiducial vector ${\mathbf{w}}:=M\mathbf{v}$. Moreover, $F^{\rm eproj}\!\left(S\right)$ is equal to the compositum of all the $F^{\rm proj}\!\left(MS\right)$ for $M\in\operatorname{EC}(d)$.
Proof.
The scalar matrices $zI$ with $\left|z\right|=1$ act trivially on projections, with $(zI\mathbf{v})(zI\mathbf{v})^{\dagger}=\mathbf{v}\mathbf{v}^{\dagger}$. The matrices $X$, $Z$, and $F$ all have entries in $\mathbb{Q}(\xi)$ (where $\xi=\xi_{d}=-e^{\frac{\pi i}{d}}$). The matrix $\sqrt{d}F$ also has entries in $\mathbb{Q}(\xi_{d})$, and $(F\mathbf{v})(F\mathbf{v})^{\dagger}=\frac{1}{d}(\sqrt{d}F\mathbf{v})(\sqrt{d}F\mathbf{v})^{\dagger}$. Therefore, writing any $M\in EC(d)$ as a product of the generators given in 3.7 shows that $(M\mathbf{v})(M\mathbf{v})^{\dagger}=M\mathbf{v}\mathbf{v}^{\dagger}M^{\dagger}$ has entries in $F^{\rm eproj}\!\left(S\right)$ (the compositum of $F^{\rm proj}\!\left(S\right)$ and $\mathbb{Q}(\xi)$). This argument shows that $F^{\rm proj}\!\left(MS\right)\subseteq F^{\rm eproj}\!\left(S\right)$, so $F^{\rm eproj}\!\left(MS\right)\subseteq F^{\rm eproj}\!\left(S\right)$. By considering $M^{-1}$, we also have $F^{\rm eproj}\!\left(S\right)=F^{\rm eproj}\!\left(M^{-1}MS\right)\subseteq F^{\rm eproj}\!\left(MS\right)$, so in fact $F^{\rm eproj}\!\left(MS\right)=F^{\rm eproj}\!\left(S\right)$.
To show that $F^{\rm eproj}\!\left(S\right)$ is equal to the compositum of all the $F^{\rm proj}\!\left(MS\right)$, it suffices to show that $\xi=-e^{-\frac{\pi i}{d}}$ occurs in the compositum of two fields $F^{\rm proj}\!\left(S\right)$ and $F^{\rm proj}\!\left(RS\right)$ where $R$ is the element of $\operatorname{C}(d)$ given in 3.7. Given a fiducial ${\mathbf{v}}$ generating $S=S([{\mathbf{v}}])$ having $v_{0}v_{1}\neq 0$, we consider the fiducial vector ${\mathbf{w}}=R{\mathbf{v}}$ of a geometrically equivalent SIC. It has entries $(R{\mathbf{v}})_{j}=\xi^{j(j+d)}v_{j}$ so $R{\mathbf{v}}_{0}=v_{0}$ and $R{\mathbf{v}}_{1}=-e^{\frac{\pi i}{d}}v_{1}$. Therefore $\Pi^{\prime}={\mathbf{w}}{\mathbf{w}}^{\dagger}$ has
$(\Pi^{\prime})_{1,0}=-e^{\frac{\pi i}{d}}v_{1}\overline{v}_{0}\in F^{\rm proj}\!\left(RS\right),$ (3.13)
while $v_{1}\overline{v}_{0}\in F^{\rm proj}\!\left(S\right)$. Now $v_{0}v_{1}\neq 0$ implies we may take their quotient in the compositum field and so get $-e^{\frac{\pi i}{d}}\in F^{\rm proj}\!\left(S\right)F^{\rm proj}\!\left(RS\right)$, as required. ∎
4. Algebraic SICs and the Galois action
We now focus on the subclass of algebraic SICs, for which the fields defined in Section 3.1 are algebraic number fields. In this paper, $\overline{\mathbb{Q}}$ is defined to be the algebraic closure of $\mathbb{Q}$ in $\mathbb{C}$, so there is a canonical choice of complex conjugation automorphism $\sigma_{c}\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$, $\sigma_{c}(z)=\overline{z}$. We show that there is a well-defined Galois action on algebraic SICs and that the property of $S$ being an algebraic SIC is preserved by the subgroup of all automorphisms $\sigma\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ that commute with $\sigma_{c}$ when restricted to the Galois closure of $F^{\rm proj}\!\left(S\right)$.
It is conjectured for dimensions $d\neq 3$ that all Weyl–Heisenberg SICs are algebraic. In Section 4.3, we define a combined geometric and Galois action on all algebraic Weyl–Heisenberg SICs and classify them into multiplets.
4.1. Algebraic SICs
We consider SICs generated by algebraic vectors.
Definition 4.1.
A complex line $\mathbb{C}{\mathbf{v}}$ generated by a nonzero vector ${\mathbf{v}}\in\mathbb{C}^{d}$ is a projective-algebraic line if all the ratios $v_{i}/v_{j}\in\overline{\mathbb{Q}}$ (for $v_{j}\neq 0$). (The definition is independent of the choice of ${\mathbf{v}}$.) A nonzero vector such that $\mathbb{C}\mathbf{v}$ is a projective-algebraic line is called a projective-algebraic vector.
An algebraic vector is some ${\mathbf{v}}\in{\overline{\mathbb{Q}}}^{d}$. For a projective-algebraic vector $\mathbf{v}$, we set its $\overline{\mathbb{Q}}$-equivalence class
$[[{{\mathbf{v}}}]]_{\operatorname{alg}}:=\{\lambda{\mathbf{v}}:\lambda\in\mathbb{C}\mbox{ and }\lambda{\mathbf{v}}\in\overline{\mathbb{Q}}^{d}\}.$ (4.1)
The next result implies that the algebraic class $[[{{\mathbf{v}}}]]_{\operatorname{alg}}$ of an algebraic complex line always includes algebraic unit vectors.
Lemma 4.2.
If ${\mathbf{v}}\in\mathbb{C}^{d}$ is a projective-algebraic vector, then there exists $\lambda\in\mathbb{C}^{\times}$ such that ${\mathbf{x}}=\lambda{\mathbf{v}}$ is algebraic and $\left|{\mathbf{x}}\right|=1$. The $\overline{\mathbb{Q}}$-equivalence class $[[{\mathbf{x}}]]_{\operatorname{alg}}$ is independent of $\lambda$.
Proof.
Let ${\mathbf{v}}$ have some $v_{i}\neq 0$ and set $\mu=\frac{1}{v_{i}}$. Then ${\mathbf{w}}:=\mu{\mathbf{v}}=\left(\frac{v_{0}}{v_{i}},\frac{v_{1}}{v_{i}},\ldots,\frac{v_{d-1}}{v_{i}}\right)\in\overline{\mathbb{Q}}^{d}$ by the projective-algebraicity of ${\mathbf{v}}$. Now the complex conjugates $\sigma_{c}\!\left(\frac{v_{k}}{v_{i}}\right)=\frac{\overline{v}_{k}}{\overline{v}_{i}}$ are algebraic numbers, hence so is $\left|\frac{v_{k}}{v_{i}}\right|^{2}$. Thus, $\left|{\mathbf{w}}\right|=\sqrt{\sum_{k}\left|\frac{v_{k}}{v_{i}}\right|^{2}}$ is a positive real algebraic number. We set $\lambda=\frac{\mu}{\left|{\mathbf{w}}\right|}$ and obtain the required ${\mathbf{x}}=\frac{{\mathbf{w}}}{\left|{\mathbf{w}}\right|}=\lambda{\mathbf{v}}$ having $\left|{\mathbf{x}}\right|=1$.
Finally, if $\lambda$ and $\lambda^{\prime}$ give two such ${\mathbf{x}},{\mathbf{x}}^{\prime}\!$, necessarily $\lambda^{\prime}/\lambda\in\overline{\mathbb{Q}}$. ∎
Definition 4.3.
A $d$-dimensional line-SIC $S=\{{\mathbf{v}}_{i}:1\leq i\leq d^{2}\}$ is algebraic if each of its $d^{2}$ vectors ${\mathbf{v}}_{i}$ are projective-algebraic.
Lemma 4.4.
Algebraic line-SICs satisfy the following properties.
  • (1)
[…]
  • (2)
    For an algebraic SIC the fields $F^{\rm vec}\!\left(S\right)$, $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$ are algebraic number fields; that is, finite algebraic extensions of $\mathbb{Q}$.
  • (3)
    A Weyl–Heisenberg line-SIC with fiducial vector ${\mathbf{v}}$ is an algebraic SIC if and only if ${\mathbf{v}}$ is a projective-algebraic vector.
Proof.
Proof of (1): This follows using 4.2.
Proof of (2): That $F^{\rm vec}\!\left(S\right)$ is an algebraic number field follows directly from the definition of algebraic. Given a line-SIC specified by $d^{2}$ projective algebraic lines $\mathbb{C}{{\mathbf{v}}_{i}}$, using (1) we immediately have $F^{\rm trip}\!\left(S\right)$ being an algebraic number field. Furthermore, one may rescale all ${\mathbf{v}}_{i}$ to be algebraic unit vectors by 4.2. Now the associated projection matrices $\Pi_{{\mathbf{v}}}={\mathbf{v}}_{i}{\mathbf{v}}_{i}^{\dagger}$ have all entries in $\overline{\mathbb{Q}}$, hence $F^{\rm proj}\!\left(S\right)$ will be an algebraic number field.
Proof of (3): If the fiducial vector ${\mathbf{v}}$ of a Weyl–Heisenberg line-SIC $S$ is projective-algebraic, then by 4.2, ${\mathbf{v}}$ may be rescaled to be an algebraic unit vector ${\mathbf{w}}=\lambda{\mathbf{v}}$. The Weyl–Heisenberg group consists of matrices with algebraic entries, hence its action on ${\mathbf{w}}$ shows that all ${\mathbf{w}}_{i}$ for $1\leq i\leq d^{2}$ are algebraic, so $S$ is an algebraic line-SIC. The converse direction is immediate. ∎
It is conjectured that for each dimension $d\neq 3$, all Weyl–Heisenberg SICs are algebraic.
Conjecture 4.5 (Weyl–Heisenberg Algebraicity Conjecture).
For each dimension $d\geq 1$, except $d=3$, all Weyl–Heisenberg SICs are projective-algebraic.
This conjecture is equivalent to 2.5, which states that for all $d\neq 3$ the real projective algebraic variety $\operatorname{WHSIC}(d)$ is a finite set. It would follow by 4.4 that, if $S$ a Weyl–Heisenberg SIC, then $F^{\rm vec}\!\left(S\right)$, $F^{\rm trip}\!\left(S\right)$, and $F^{\rm proj}\!\left(S\right)$ are algebraic number fields.
4.2. Galois actions on algebraic SICs
There is a well-defined action of $\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ on algebraic complex lines which maps them to other algebraic complex lines. Given $\sigma\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ and ${\mathbf{v}}\in\overline{\mathbb{Q}}^{d}$ we send ${\mathbf{v}}\mathbb{C}\mapsto\sigma({\mathbf{v}})\mathbb{C}$. This action is well-defined on algebraic complex lines, since if $\lambda\in\overline{\mathbb{Q}}^{\times}$, then ${\mathbf{v}}\mathbb{C}=\lambda{\mathbf{v}}\mathbb{C}$ while
$\sigma(\lambda{\mathbf{v}})\mathbb{C}=\sigma(\lambda)\sigma({\mathbf{v}})\mathbb{C}=\sigma({\mathbf{v}})\mathbb{C}.$ (4.2)
Under this Galois action, the image under $\sigma$ of $d^{2}$ algebraic complex lines will be $d^{2}$ algebraic complex lines. If these algebraic lines form a line-SIC, then the image set of lines can be either the same SIC, a different SIC, or not a SIC at all.
For a particular SIC, there is a restricted set of Galois automorphisms (which may depend on the line-SIC) that map it to another line-SIC.
Definition 4.6.
Let $\sigma_{c}\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ denote complex conjugation and $E$ an algebraic Galois extension of $\mathbb{Q}$. Let $C(\sigma_{c}\!\mid\!E)$ be the centralizer of $\sigma_{c}$ restricted to $\operatorname{Gal}(E/\mathbb{Q})$, that is,
$C(\sigma_{c}\!\mid\!E):=\{\sigma\in\operatorname{Gal}(E/\mathbb{Q}):\,\sigma\sigma_{c}(\alpha)=\sigma_{c}\sigma(\alpha)\mbox{ for all }\alpha\in E\}.$ (4.3)
It is a subgroup of $\operatorname{Gal}(E/\mathbb{Q})$.
For a line-SIC $S$, we will show there is a Galois field $E$ (depending on $S$) for which $C(\sigma_{c}\!\mid\!E)$ is such a restricted set. However, neither the field $F^{\rm proj}\!\left(S\right)$ nor $F^{\rm eproj}\!\left(S\right)$ is known to always be a Galois extension of $\mathbb{Q}$, although there is no known counterexample for $d\neq 3$. Thus, we give a name to the Galois closure of $F^{\rm eproj}\!\left(S\right)$.
Definition 4.7.
Given an algebraic line-SIC $S$, let $F^{\rm Gproj}\!\left(S\right)$ denote the normal closure of $F^{\rm eproj}\!\left(S\right)$ over $\mathbb{Q}$, and call it the Galois-projection SIC field of $S$.
The next proposition shows that $C(\sigma_{c}\!\mid\!F^{\rm Gproj}\!\left(S\right))$ is a group of Galois automorphisms that preserve the property of being an algebraic line-SIC. For $d\geq 1$, the group $C(\sigma_{c}\!\mid\!F^{\rm Gproj}\!\left(S\right))$ is always nontrivial because it contains complex conjugation, which acts nontrivially because $F^{\rm proj}\!\left(S\right)$ cannot be contained in $\mathbb{R}$. (A set of real equiangular lines in $\mathbb{R}^{d}$ has cardinality at most $\frac{d(d+1)}{2}<d^{2}$.)
Proposition 4.8.
Let $S$ be an algebraic line-SIC. If $\sigma\in C(\sigma_{c}\!\mid\!F^{\rm Gproj}\!\left(S\right))$, then $\sigma(S)$ is also an algebraic line-SIC, and $F^{\rm Gproj}\!\left(\sigma(S)\right)=F^{\rm Gproj}\!\left(S\right)$.
Proof.
Write the line-SIC and (projection) SIC as
$S$ =$=\left\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{d^{2}}\right\},$ (4.4)
${\mathbf{\Pi}}(S)$ =$=\left\{\tfrac{1}{d}\Pi_{1},\tfrac{1}{d}\Pi_{2},\ldots,\tfrac{1}{d}\Pi_{d^{2}}\right\}.$ (4.5)
After rescaling if needed, all the entries of the $\mathbf{v}_{i}$ lie in $F^{\rm vec}\!\left(S\right)$, hence they lie in $F^{\rm proj}\!\left(S\right)$, so they lie in $F^{\rm Gproj}\!\left(S\right)$. The entries of the $\Pi_{i}$ also lie in $F^{\rm proj}\!\left(S\right)$, hence $F^{\rm Gproj}\!\left(S\right)$. Also,
$\sigma(\Pi_{i})=\sigma(\mathbf{v}_{i}\sigma_{c}(\mathbf{v}_{i})^{\top})=\sigma(\mathbf{v}_{i})\sigma(\sigma_{c}(\mathbf{v}_{i})^{\top})=\sigma(\mathbf{v}_{i})\sigma_{c}(\sigma(\mathbf{v}_{i})^{\top})$ (4.6)
is the Hermitian projection onto $\sigma(\mathbf{v}_{i})$, so
${\mathbf{\Pi}}(\sigma(S))=\left\{\tfrac{1}{d}\sigma(\Pi_{1}),\tfrac{1}{d}\sigma(\Pi_{2}),\ldots,\tfrac{1}{d}\sigma(\Pi_{d^{2}})\right\}.$ (4.7)
Moreover,
$\operatorname{Tr}(\sigma(\Pi_{i})\sigma(\Pi_{j}))=\sigma(\operatorname{Tr}(\Pi_{i}\Pi_{j}))=\sigma\!\left(\tfrac{\delta_{ij}d+1}{d+1}\right)=\tfrac{\delta_{ij}d+1}{d+1},$ (4.8)
so ${\mathbf{\Pi}}(\sigma(S))$ is a SIC. Finally, $F^{\rm Gproj}\!\left(\sigma(S)\right)\subseteq F^{\rm Gproj}\!\left(S\right)$, because the entries of $\Pi_{i}$ lie in $F^{\rm Gproj}\!\left(S\right)$, and $F^{\rm Gproj}\!\left(S\right)$ is a Galois field fixed by $\sigma$. It follows by symmetry that $F^{\rm Gproj}\!\left(S\right)=F^{\rm Gproj}\!\left(\sigma^{-1}(\sigma(S))\right)\subseteq F^{\rm Gproj}\!\left(\sigma(S)\right)$ and thus $F^{\rm Gproj}\!\left(\sigma(S)\right)=F^{\rm Gproj}\!\left(S\right)$. ∎
4.3. Galois multiplets and super-equivalence classes of Weyl–Heisenberg SICs
We now assemble $\operatorname{PEC}(d)$-orbits of Weyl–Heisenberg SICs in dimension $d$ into multiplets under the Galois action by the subgroup $C(\sigma_{c}\!\mid\!F^{\rm eproj}\!\left(S\right))$ of $\operatorname{Gal}(F^{\rm eproj}\!\left(S\right)/\mathbb{Q})$ that commutes with complex conjugation. The coarser equivalence classes obtained under the combined geometric and Galois actions (i.e., unions of the orbits in a multiplet) will called Galois super-equivalence classes.
The conjecture of Appleby, Flammia, McConnell, and Yard [10, Conj. 2] implies that many of the fields we have been studying are equivalent Galois extensions of $\mathbb{Q}$.
Lemma 4.9.
Assume 1.1 holds. Then in every dimension $d\geq 4$, for each Weyl–Heisenberg line-SIC $S$, the field $F^{\rm vec}\!\left(S\right)$ is a Galois extension of $\mathbb{Q}$, and
$F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)=F^{\rm eproj}\!\left(S\right)=F^{\rm Gproj}\!\left(S\right).$ (4.9)
Proof.
By assertion (2) of 1.1 $F^{\rm vec}\!\left(S\right)$ is a Galois extension of $\mathbb{Q}$. Consequently, $F^{\rm vec}\!\left(S\right)$ is preserved by complex conjugation, so $F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)$ holds by 3.5(1).
By assertion (2) of 1.1 $F^{\rm vec}\!\left(S\right)$ always contains the ray class field $H_{d^{\prime}\infty_{1}\infty_{2}}$ of $\mathbb{Q}(\Delta_{d})$. This class field must contain the class field of $\mathbb{Q}$ with conductor $d^{\prime}\infty$, which is $\mathbb{Q}(\zeta_{d^{\prime}})=\mathbb{Q}(\xi_{d})$. Therefore, $F^{\rm vec}\!\left(S\right)$ contains $F^{\rm proj}\!\left(S\right)(\xi_{d})=F^{\rm eproj}\!\left(S\right)$. Since $F^{\rm vec}\!\left(S\right)\subseteq F^{\rm proj}\!\left(S\right)\subseteq F^{\rm eproj}\!\left(S\right)$, the equality of those three fields follows. So $F^{\rm eproj}\!\left(S\right)$ is Galois and thus also equal to $F^{\rm Gproj}\!\left(S\right)$. ∎
We extract the following conjecture, of which all but the claim about $F^{\rm trip}\!\left(S\right)$ being Galois is already a consequence of 1.1.
Conjecture 4.10.
Fix a positive integer $d\neq 3$. For any $d$-dimensional Weyl–Heisenberg line SIC $S$, its $d^{2}$ vectors are always projective-algebraic. The associated fields $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$ (and thus also $F^{\rm eproj}\!\left(S\right)$) are Galois extensions of $\mathbb{Q}$.
The evidence for a Galois action on Weyl–Heisenberg SICs was extensively detailed by Appleby, Yadsan-Appleby, and Zauner [4] in 2013.
The group $\operatorname{Gal}(\overline{\mathbb{Q}})/\mathbb{Q})$ acts on the elements of Weyl–Heisenberg group $\operatorname{WH}(d)$ coordinatewise on its matrices. Any automorphism $\sigma:\overline{\mathbb{Q}}\to\overline{\mathbb{Q}}$, sending a matrix $H\mapsto\sigma(H)$, acts as a homomorphism $\sigma(H_{1}H_{2})=\sigma(H_{1})\sigma(H_{2})$. This Galois action fixes the group $\operatorname{WH}(d)$ as a set, with any automorphism permuting its elements, since every Galois conjugate $\sigma(H)$ of a group element $H$ is also an element of $\operatorname{WH}(d)$.
By 3.8(2), $F^{\rm eproj}\!\left(S\right)$ is an invariant of the $\operatorname{PEC}(d)$ orbit of the Weyl–Heisenberg SIC $S$. If we assume 4.10, then $F^{\rm eproj}\!\left(S\right)$ is preserved by the Galois action by the subgroup $C(\sigma_{c}\!\mid\!F^{\rm eproj}\!\left(S\right))$ of $\operatorname{Gal}(F^{\rm eproj}\!\left(S\right)/\mathbb{Q})$. In other words, the extended projection field is invariant under both the geometric and the Galois actions.
We make the following definition.
Definition 4.11.
Assume 4.10 holds. For $d\geq 4$ the Galois multiplet of a $\operatorname{PEC}(d)$-equivalence class $[S]$ of Weyl–Heisenberg line-SICs is
$[[S]]=\{[\sigma(S)]:\sigma\in C(\sigma_{c}\!\mid\!F^{\rm eproj}\!\left(S\right))\},$ (4.10)
that is, the set of Galois equivalent orbits. We also define the equivalence relation $\approx_{GPEC}$, Galois super-equivalence, by
$S_{1}\approx_{GPEC}S_{2}\iff[[S_{1}]]=[[S_{2}]],$ (4.11)
and term the equivalence class of a SIC $S$ under $\approx_{GPEC}$ (that is, the union of the elements of $[[S]]$) the Galois super-equivalance class of $S$.
Remark 4.12.
Assuming 1.1 holds, this definition coincides with that of “Galois multiplet” in [9].
We summarize the properties noted above in the following proposition.
Proposition 4.13.
Assume 4.10 holds. Then, the set $\operatorname{WHSIC}(d)$ is divided into a finite number of Galois super-equivalence classes under the equivalence relation $\approx_{GPEC}$. Each Galois super-equivalence class is a finite union of (geometric) $\operatorname{PEC}(d)$-equivalence classes.
Galois multiplets are studied in Sections 7 and 8 using data and more refined conjectures. The multiplets in a fixed dimension often have different sizes.
5. A conjecture on counting geometric equivalence classes of Weyl–Heisenberg SICs
In this section, we formulate a conjecture relating the number of SICs in dimensions $d\geq 4$ to class numbers of certain real quadratic orders. We present numerical evidence.
5.1. Class groups and class monoids of orders
Let $K$ be a number field. An order in $K$ is any subring $\mathcal{O}$ including $1$ that is finitely additively generated and whose rational span $\mathbb{Q}\mathcal{O}=K$. There is a unique maximal order $\mathcal{O}_{K}$ of $K$ that contains all the others; $\mathcal{O}_{K}$ coincides with the set of algebraic integers in $K$. The maximal order is a Dedekind domain, while every non-maximal order is not a Dedekind domain. In particular, a non-maximal order will have non-invertible nonzero ideals. However, the set of invertible ideals of a non-maximal order behaves much like set of nonzero ideals in a Dedekind domain.
Recall that a fractional ideal of a (commutative) integral domain $\mathcal{O}$ is a finitely generated $\mathcal{O}$-module $\mathfrak{m}$ contained in its quotient field $K$. A fractional ideal $\mathfrak{a}$ is invertible if there exists a fractional ideal $\mathfrak{b}$ with $\mathfrak{a}\mathfrak{b}=\mathcal{O}$. Dedekind domains are characterized by the property that all nonzero fractional ideals are invertible.
The class group of an order, defined in terms of invertible fractional ideals, generalizes the usual notion of the class group of (the maximal order of) a number field.
Definition 5.1.
The class group (also called the Picard group or ring class group) of an order $\mathcal{O}$ is
$\operatorname{Cl}(\mathcal{O}):=\frac{\{\mbox{invertible fractional ideals of }\mathcal{O}\}}{\{\mbox{principal fractional ideals of }\mathcal{O}\}}.$ (5.1)
If one instead considers all nonzero ideals, one obtains a monoid (that is, a semigroup with identity) rather than a group.
Definition 5.2.
The class monoid of an order $\mathcal{O}$ is
$\operatorname{Clm}(\mathcal{O}):=\frac{\{\mbox{nonzero fractional ideals of }\mathcal{O}\}}{\{\mbox{principal fractional ideals of }\mathcal{O}\}}.$ (5.2)
For a more in-depth discussion of the nuances of orders, see [33] and [35, Ch. I, Sec. 12].
If $K$ is a quadratic field of discriminant $\Delta_{0}$, it has a unique suborder of discriminant $f^{2}\Delta_{0}$ for every positive integer $f$. This order is given explicitly as
$\mathcal{O}_{\!\Delta}:=\mathbb{Z}\!\left[\frac{\Delta+\sqrt{\Delta}}{2}\right]=\mathbb{Z}+\frac{\Delta+\sqrt{\Delta}}{2}\mathbb{Z}=\mathbb{Z}+f\frac{\Delta_{0}+\sqrt{\Delta_{0}}}{2}\mathbb{Z}.$ (5.3)
For the SIC problem, we are specifically interested in the case when $\Delta=\Delta_{d}=(d+1)(d-3)$ for some integer $d$. In this case, the class monoid $\operatorname{Clm}(\mathcal{O}_{\!\Delta_{d}})$ may be understood through bijections with other sets of number-theoretic interest.
Theorem 5.3.
Suppose $d\in\mathbb{Z}\setminus\{-1,3\}$, and let $\Delta=(d+1)(d-3)$. Write $\Delta_{d}=f^{2}\Delta_{0}$ for some $f\in\mathbb{N}$ and a fundamental discriminant $\Delta_{0}$. Then there are explicit, canonical bijections between the sets
  • (0)
    $\operatorname{Clm}(\mathcal{O}_{\Delta})$;
  • (1)
    the disjoint union $\bigsqcup_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$;
  • (2)
    the set $\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})$, where $\mathcal{Q}(\Delta)$ is the set of binary quadratic forms of discriminant $\Delta$, and ${A}=\left(\begin{smallmatrix}s&t\\ u&v\end{smallmatrix}\right)\in\mathbf{GL}_{2}(\mathbb{Z})$ acts on the right by “twisted precomposition” $Q^{{A}}(x,y)=(\det{{A}})Q(sx+ty,ux+vy)$; and
  • (3)
    the set $\mathcal{C}_{d-1}=\{\{{R}^{-1}{A}{R}:{R}\in\mathbf{GL}_{2}(\mathbb{Z})\}:{A}\in\mathbf{SL}_{2}(\mathbb{Z}),\operatorname{Tr}({A})=d-1\}$ of $\mathbf{GL}_{2}(\mathbb{Z})$-conjugacy class of elements of $\mathbf{SL}_{2}(\mathbb{Z})$ of trace $d-1$.
Proof.
The bijection between (0) and (1) holds for quadratic orders in general. Specicially, by [33, Prop. A.12 and Prop. A.4], $\operatorname{Clm}(\mathcal{O}_{\Delta})=\operatorname{Clm}^{\flat}(\mathcal{O}_{\Delta})\cong\bigsqcup_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$.
The bijection between (0) and (2) also holds for quadratic orders in general (and with appropriate modifications, nondegenerate quadratic rings) and is known as twisted Gauss composition. This is proven as [47, Thm. 1.2]; for an introduction to Gauss composition, see [16, Ch. 5].
We now exhibit a map $\phi:\mathcal{Q}(\Delta)\to\mathbf{GL}_{2}(\mathbb{Z})$ that induces a bijection between (2) and (3). For a quadratic form $Q(x,y)=ax^{2}+bxy+cy^{2}$ of discriminant $\Delta=(d+1)(d-3)$, define $\phi(Q):=\left(\begin{smallmatrix}\frac{1}{2}(b+d-1)&-a\\ c&\frac{1}{2}(-b+d-1)\end{smallmatrix}\right)$. Note that $\det(\phi(Q))=\frac{1}{4}(-b^{2}+(d-1)^{2})+ac=\frac{1}{4}(-\Delta+(d-1)^{2})=1$ and $\operatorname{Tr}(\phi(Q))=d-1$. It is straightforward to check that $\phi(Q^{R})={R}^{-1}\phi(Q){R}$ for ${R}\in\mathbf{GL}_{2}(\mathbb{Z})$. Thus, $\phi$ defines a map $\overline{\phi}:\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})\to\mathcal{C}_{d-1}$. In the other direction, for ${A}=\left(\begin{smallmatrix}s&t\\ u&v\end{smallmatrix}\right)\in SL_{2}(\mathbb{Z})$, define the quadratic form $(\psi({A}))(x,y):=-tx^{2}+(s-v)xy+uy^{2}$. It is straightforward to check that $\operatorname{disc}(\psi({A}))=\Delta$ and $\psi({R}^{-1}{A}{R})=\psi({A})^{R}$, so $\psi$ defines a map $\overline{\psi}:\mathcal{C}_{d-1}\to\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})$. A direct calculation shows $\overline{\phi}\circ\overline{\psi}=\operatorname{id}$ and $\overline{\psi}\circ\overline{\phi}=\operatorname{id}$. ∎
5.2. Statement of the Geometric Count Conjecture
The following conjecture counts the number of geometric equivalence classes of Weyl–Heisenberg SICs in dimension $d$.
Conjecture 5.4 (Geometric Count Conjecture).
Fix a positive integer $d\neq 3$, and let $\Delta=(d+1)(d-3)$ and $\mathcal{O}_{\Delta}$ be the quadratic order of discriminant $\Delta$. Then, the number of geometric equivalence classes $[S]$ of Weyl–Heisenberg covariant SICs is
$\left|{\rm WHSIC}(d)/\operatorname{PEC}(d)\right|=\left|\operatorname{Clm}(\mathcal{O}_{\Delta})\right|.$ (5.4)
By 5.3, the quantity $\left|\operatorname{Clm}(\mathcal{O}_{\Delta})\right|$ is equal to
  • (1)
    the sum of (ring) class numbers $\sum_{f^{\prime}|f}h_{(f^{\prime})^{2}\Delta_{0}}$, where $h_{(f^{\prime})^{2}\Delta_{0}}=\left|\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\right|$;
  • (2)
    $\left|\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})\right|$, the number of twisted $\mathbf{GL}_{2}(\mathbb{Z})$-classes of binary quadratic forms of discriminant $\Delta$;
  • (3)
    $\left|\mathcal{C}_{d-1}\right|$, the number of $\mathbf{GL}_{2}(\mathbb{Z})$-conjugacy class of elements of $\mathbf{SL}_{2}(\mathbb{Z})$ of trace $d-1$.
5.3. Verification of cases of the Geometric Count Conjecture
We have verified that the two sets in the Geometric Count conjecture have the same size for dimensions $d\leq 90$, $d\neq 3$ under the assumption that the published lists of Weyl–Heisenberg SICs in these dimensions are complete.
Scott [38] gives a putatively complete list of numerical SICs in all dimensions $d\leq 90$, building on earlier work of Scott and Grassl [39]. Scott and Grassl find SIC fiducials by numerically minimizing a frame potential—a function whose global minima are Weyl–Heisenberg SIC fiducials. By repeating their search with a large number of seed vectors, they establish a high degree of numerical confidence in their results but do not rigorously rule out the existence of other SICs.
The dimensions $d\in\{1,2\}$ are special cases—the only dimensions for which $\Delta<0$, or equivalently, the conjugacy classes in $\mathcal{C}_{d-1}$ are elliptic. For $d=1$, there is exactly one SIC ($\mathbb{C}^{1}$ as a line in $\mathbb{C}^{1}$), and it is (trivially) Weyl–Heisenberg covariant. The associated order $\mathcal{O}_{-4}=\mathbb{Z}[i]$ is a maximal order of class number one, so the conjecture is true. For $d=2$, there is again a unique SIC up to unitary transformations (a regular tetrahedron on the Riemann sphere $\mathbb{P}^{1}(\mathbb{C})$), and it may be chosen to be Weyl–Heisenberg covariant. The associated order $\mathcal{O}_{-3}=\mathbb{Z}[e^{2\pi i/3}]$ is a maximal order of class number one, so the conjecture is true.
5.4 omits the case $d=3$, where $\Delta=0$, where associated field $K_{\!\Delta}=\mathbb{Q}$ has $\mathcal{O}_{\!\Delta}=\operatorname{Z}$. In this case, there is a continuous one-parameter family of Weyl–Heisenberg covariant SICs [39]. For $d=3$ the classification (3) above is the set $\mathcal{C}_{2}$ of parabolic conjugacy classes of $\mathbf{SL}_{2}(\mathbb{Z})$, which is infinite.
For $4\leq d\leq 90$ we computed $\left|\operatorname{Clm}(\mathcal{O}_{\Delta})\right|$ using Magma [17] by computing the equivalent quantity $\left|\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})\right|$. Write $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ for a fundamental discriminant $\Delta_{0}$. For each $f^{\prime}\mid f$, we computed the set of reduced primitive binary quadratic forms of discriminant $(f^{\prime})^{2}\Delta_{0}$ using the Magma function ReducedForms(), written by David Kohel. This function produces a list of $\mathbf{SL}_{2}(\mathbb{Z})$-classes of primitive binary quadratic forms in which a twisted $\mathbf{GL}_{2}(\mathbb{Z})$-class is always represented either by a single form $ax^{2}+bxy+cy^{2}$ with $a>0$ or a pair of forms $\{ax^{2}+bxy+cy^{2},-ax^{2}+bxy-cy^{2}\}$. Removing forms with $a<0$ produced a list containing one representative for each primitive twisted $\mathbf{GL}_{2}(\mathbb{Z})$-class. Multiplying by $f/f^{\prime}$ and joining the lists gave a list of representatives of all (not necessarily primitive) classes in $\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})$.
We observe that for each $d$ ($4\leq d\leq 90$) that the size of
$\operatorname{Clm}(\mathcal{O}_{\Delta})=\sum_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$ (5.5)
is equal to the sum of the corresponding values in the “#” column of Scott’s table beginning on page 3 of [38], which give the number of elements of ${\rm WHSIC}(d)/\operatorname{PEC}(d)$. This is agreement is shown in Table 1. The table also includes the sizes of the ring class groups $\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$ (equivalently, the number of twisted $\mathbf{GL}_{2}(\mathbb{Z})$-classes of primitive binary quadratic forms with discriminant $(f^{\prime})^{2}\Delta_{0}$).
$d$ $\Delta_{0}$ $f^{\prime}$ $h_{(f^{\prime})^{2}\Delta_{0}}$#\##
$1$ $-4$ $1$ $1$ $1$
$2$ $-3$ $1$ $1$ $1$
$3$ 00 $\ast$ $\ast$ $\infty$
$4$ $5$ $1$ $1$ $1$
$5$ $12$ $1$ $1$ $1$
$6$ $21$ $1$ $1$ $1$
$7$ $8$ $1,2$ $1,1$ $2$
$8$ $5$ $1,3$ $1,1$ $2$
$9$ $60$ $1$ $2$ $2$
$10$ $77$ $1$ $1$ $1$
$11$ $24$ $1,2$ $1,2$ $3$
$12$ $31$,3131,31 , 3 $1,1$ $2$
$13$ $140$ $1$ $2$ $2$
$14$ $165$ $1$ $2$ $2$
$15$ $21$,2,41241,2,41 , 2 , 4 $1,1,2$ $4$
$16$ $221$ $1$ $2$ $2$
$17$ $28$ $1,3$ $1,2$ $3$
$18$ $285$ $1$ $2$ $2$
$19$ $5$ $1,2,4,8$ $1,1,1,2$ $5$
$20$ $357$ $1$ $2$ $2$
$21$ $44$ $1,3$ $1,4$ $5$
$22$ $437$ $1$ $1$ $1$
$23$ $201$,2121,21 , 2 $2,4$ $6$
$24$ $21$ $1,5$ $1,2$ $3$
$25$ $572$ $1$ $2$ $2$
$26$ $69$ $1,3$ $1,3$ $4$
$27$ $681$,2121,21 , 2 $2,4$ $6$
$28$ $29$ $1,5$ $1,2$ $3$
$29$ $780$ $1$ $4$ $4$
$30$ $93$ $1,3$ $1,3$ $4$
$31$ $56$ $1,2,4$ $1,2,4$ $7$
$32$ $957$ $1$ $2$ $2$
$34$ $1020$ $1$ $4$ $4$
$34$ $1085$ $1$ $2$ $2$
$35$ $8$ $1,2,3,4,6,12$ $1,1,1,1,2,4$ $10$
$36$ $1221$ $1$ $4$ $4$
$37$ $1292$ $1$ $4$ $4$
$38$ $1365$ $1$ $4$ $4$
$39$ $40$ $1,2,3,6$ $2,2,2,4$ $10$
$40$ $1517$ $1$ $2$ $2$
$41$ $1596$ $1$ $8$ $8$
$42$ $1677$ $1$ $4$ $4$
$43$ $440$ $1,2$ $2,4$ $6$
$44$ $205$ $1,3$ $2,4$ $6$
$45$ $1932$ $1$ $4$ $4$
 $d$ $\Delta_{0}$ $f^{\prime}$ $h_{(f^{\prime})^{2}\Delta_{0}}$#\## $46$ $2021$ $1$ $3$ $3$ $47$ $33$ $1,2,4,8$ $1,1,2,4$ $8$ $48$ $5$ $1,3,7,21$ $1,1,1,4$ $7$ $49$ $92$ $1,5$ $1,6$ $7$ $50$ $2397$ $1$ $2$ $2$ $51$ $561$,2,41241,2,41 , 2 , 4 $2,4,8$ $14$ $52$ $53$ $1,7$ $1,3$ $4$ $53$ $21$,3,$5,1$ 3 , 5 , 15 $1,1,2,6$ $10$ $54$ $2805$ $1$ $4$ $4$ $55$ $728$ $1,2$ $2,4$ $6$ $56$ $3021$ $1$ $6$ $6$ $57$ $348$ $1,3$ $2,6$ $8$ $58$ $3245$ $1$ $4$ $4$ $59$ $840$ $1,2$ $4,8$ $12$ $60$ $3477$ $1$ $4$ $4$ $61$ $3596$ $1$ $6$ $6$ $62$ $413$ $1,3$ $1,4$ $5$ $63$ $60$ $1,2,4,8$ $2,2,4,8$ $16$ $64$ $3965$ $1$ $4$ $4$ $65$ $4092$ $1$ $8$ $8$ $66$ $469$ $1,3$ $3,6$ $9$ $67$ $71$,2,4,8,161248161,2,4,8,161 , 2 , 4 , 8 , 16 $1,1,1,2,4$ $9$ $68$ $4485$ $1$ $4$ $4$ $69$ $4620$ $1$ $8$ $8$ $70$ $4757$ $1$ $5$ $5$ $71$ $361$,2,3,612361,2,3,61 , 2 , 3 , 6 $2,4,4,8$ $18$ $72$ $5037$ $1$ $4$ $4$ $73$ $5180$ $1$ $4$ $4$ $74$ $213$ $1,5$ $1,6$ $7$ $75$ $521$,2,3,612361,2,3,61 , 2 , 3 , 6 $1,2,4,8$ $15$ $76$ $5621$ $1$ $6$ $6$ $77$ $5772$ $1$ $8$ $8$ $78$ $237$ $1,5$ $1,6$ $7$ $79$ $380$ $1,2,4$ $2,4,8$ $14$ $80$ $77$ $1,3,9$ $1,2,6$ $9$ $81$ $6396$ $1$ $12$ $12$ $82$ $6557$ $1$ $3$ $3$ $83$ $051$,2,4,812481,2,4,81 , 2 , 4 , 8 $2,2,4,8$ $16$ $84$ $85$ $1,3,9$ $2,2,6$ $10$ $85$ $7052$ $1$ $4$ $4$ $86$ $7221$ $1$ $10$ $10$ $87$ $8481$,2121,21 , 2 $4,8$ $12$ $88$ $7565$ $1$ $4$ $4$ $89$ $860$ $1,3$ $2,8$ $10$ $90$ $7917$ $1$ $4$ $4$
Table 1. Comparison of ring class numbers and counts of SICs for $d\leq 90$. Here $(d+1)(d-3)=f^{2}\Delta_{0}$ for a fund. disc. $\Delta_{0}$, $f^{\prime}$ runs over divisors of $f$, $h_{(f^{\prime})^{2}\Delta_{0}}=\left|\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\right|$, and ##\## counts W–H SICs of Scott [38] in dim. $d$.
6. Theorems on class monoids of real quadratic fields
In this section, we prove rigorous results about the behavior of $\operatorname{Clm}(\mathcal{O})$ for general orders. These results have conditional consequences for SICs via 5.4.
6.1. Growth of the number of SICs in dimension $d$
Going beyond Zauner’s conjecture, one would like to know not only that SICs exist in every dimension $d$, but also how many there are. The connection between SICs and class numbers gives us a precise prediction for the number of SICs up to equivalence (excluding any SICs that are not Weyl–Heisenberg covariant, to which our conjectures are agnostic). In this section, we use known results on the growth of class numbers to derive conditional results on the number of SICs as $d\to\infty$.
The first growth result characterizes those dimensions in which our conjectures predict a unique Weyl–Heisenberg SIC, using a theorem of Byeon, Kim, and Lee [13].
Proposition 6.1.
One has $\left|\operatorname{Clm}(\mathcal{O}_{\!\Delta_{d}})\right|>1$ for all $d>22$. The values of $d\neq 3$ with $\left|\operatorname{Clm}(\mathcal{O}_{\!\Delta_{d}})\right|=1$ are $d\in\{1,2,4,5,6,10,22\}$.
Proof.
Set $\Delta=\Delta_{d}=(d+1)(d-3)$ and $s(d):=\left|\operatorname{Clm}(\mathcal{O}_{\!\Delta})\right|$; we are trying to determine when $s(d)=1$. By 5.4,
$s(d)=\sum_{f^{\prime}|f}h_{(f^{\prime})^{2}\Delta_{0}}.$ (6.1)
where $\Delta=f^{2}\Delta_{0}$ and $\Delta_{0}$ is fundamental.
If $f>1$, then $s(d)>1$.
If $f=1$, then $\Delta=\Delta_{0}$ is fundamental. Write $\Delta=(d-1)^{2}-4=n^{2}-4$, where $n:=d-1$. Byeon, Kim, and Lee solved the class number 1 problem for real quadratic fields with fundamental discriminant of the form $n^{2}-4$, a result previously known as Mollin’s conjecture [13, Thm 1.2]; whenever $n>21$ (so $d>22$), $s(d)=h_{\Delta}>1$. The list of dimensions where $s(d)=1$ follows from a finite calculation of class numbers. ∎
Corollary 6.2.
Assuming 5.4, there is more than one $\operatorname{PEC}(d)$-class of Weyl–Heisenberg SIC in dimension $d$ for all $d>22$. The dimensions $d$ with a unique $\operatorname{PEC}(d)$-class of Weyl–Heisenberg SIC are $d\in\{1,2,4,5,6,10,22\}$.
Proof.
The corollary follows from 6.1. ∎
The second growth result is an asymptotic formula for the size of the ray class monoid of $\mathcal{O}_{\!\Delta_{d}}$, which is conjectured to be the equal to the number of extended Clifford orbits of SICs in dimension $d$. The implied constant in this formula is ineffective because it ultimately relies on a Siegel bound involving a possible Siegel zero of a Dirichlet $L$-function. We are grateful to Steven Flammia for independently finding numerical evidence for this asymptotic [23].
Theorem 6.3.
As $d\to\infty$, and $\Delta_{d}=(d+1)(d-3)$ the size of the class monoid of $\mathcal{O}_{\!\Delta_{d}}$ obeys the asymptotic formula
$\log\left|\operatorname{Clm}(\mathcal{O}_{\!\Delta_{d}})\right|=\log d+o(\log d).$ (6.2)
Proof.
The proof uses the Brauer–Siegel theorem, which is an (ineffective) estimate on the growth of the product of the class number and the regulator. The regulator is always negligibly small in our regime. The subtlety of the proof is entirely in accounting for the class numbers of non-maximal orders, which are not included in the Brauer–Siegel theorem.
As in the proof of 6.1, set $s(d):=\left|\operatorname{Clm}(\mathcal{O}_{\Delta_{d}})\right|$. By 5.3,
$s(d)=\sum_{f^{\prime}|f}h_{(f^{\prime})^{2}\Delta_{0}}$ (6.3)
where $\Delta=\Delta_{d}=(d+1)(d-3)=f^{2}\Delta_{0}$ and $\Delta_{0}$ is fundamental. Also, $h_{(f^{\prime})^{2}\Delta_{0}}\leq h_{\Delta}$ for all $f^{\prime}\mid f$. It follows that
$h_{\Delta}\leq s(d)\leq\sigma_{0}(f)h_{\Delta},$ (6.4)
where $\sigma_{0}(f)$ is the number of positive integer divisors of $f$. Moreover,
$\log\sigma_{0}(f)\ll\frac{\log f}{\log\log f}\ll\frac{\log d}{\log\log d}$ (6.5)
by [34, Thm. 2.11]. Thus,
$\log s(d)=\log h_{\Delta}+O\!\left(\frac{\log d}{\log\log d}\right).$ (6.6)
We then use the formula for the class number of an order given as [35, Thm. 12.12]:
$h_{\Delta}=\frac{h_{\!\Delta_{0}}}{[\mathcal{O}_{\!\Delta_{0}}^{\times}:\mathcal{O}_{\!\Delta}^{\times}]}\frac{\left|\left(\mathcal{O}_{\!\Delta_{0}}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}{\left|\left(\mathcal{O}_{\Delta}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}.$ (6.7)
We will estimate the size of each factor in this formula. For any real nonsquare discriminant $\Delta^{\prime}$, let $\varepsilon_{\!\Delta^{\prime}}$ be the smallest unit of $\mathcal{O}_{\!\Delta^{\prime}}$ such that $\varepsilon_{\!\Delta^{\prime}}>1$, and let $\varepsilon_{\!\Delta^{\prime}}^{+}$ be the smallest such totally positive unit. The Brauer–Siegel theorem [12]—specifically, Siegel’s part, concerning real quadratic fields [40]— gives the asymptotic formula
$\frac{\log\!\left(h_{\!\Delta_{0}}\operatorname{Reg}(\mathbb{Q}(\Delta_{0}))\right)}{\log{\Delta_{0}}}\sim 1\mbox{ as }\Delta_{0}\to\infty.$ (6.8)
Moreover,
$[\mathcal{O}_{\!\Delta_{0}}^{\times}:\mathcal{O}_{\!\Delta}^{\times}]=\frac{\log\varepsilon_{\!\Delta}}{\log\varepsilon_{\!\Delta_{0}}}=\frac{\log\varepsilon_{\!\Delta}}{\log\operatorname{Reg}(\mathbb{Q}(\Delta_{0}))}.$ (6.9)
Thus, (6.8) may be rewritten as
$\log\!\left(\frac{h_{\!\Delta_{0}}}{[\mathcal{O}_{\!\Delta_{0}}^{\times}:\mathcal{O}_{\!\Delta}^{\times}]}\log\varepsilon_{\Delta}\right)\sim\log{\Delta_{0}}\mbox{ as }\Delta_{0}\to\infty.$ (6.10)
Fix some $\epsilon>0$. Then, there exists some natural number $N$ such that, whenever $\Delta_{0}>N$,
$(1-\epsilon)\log\Delta_{0}-\log\log\varepsilon_{\Delta}<\log\!\left(\frac{h_{\!\Delta_{0}}}{[\mathcal{O}_{\!\Delta_{0}}^{\times}:\mathcal{O}_{\!\Delta}^{\times}]}\right)<(1+\epsilon)\log\Delta_{0}-\log\log\varepsilon_{\Delta}.$ (6.11)
For $\Delta_{0}\leq N$, there is some constant $B$ such that
$B-\log\log\varepsilon_{\Delta}\leq\log\!\left(\frac{h_{\!\Delta_{0}}}{[\mathcal{O}_{\!\Delta_{0}}^{\times}:\mathcal{O}_{\!\Delta}^{\times}]}\right)\leq B-\log\log\varepsilon_{\Delta}.$ (6.12)
We may bound $\left|\left(\mathcal{O}_{\!\Delta_{0}}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|$ above by $f^{2}$ and below by
$\left|\left(\mathcal{O}_{\!\Delta_{0}}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|$ $=f^{2}\prod_{\mathfrak{p}|f}\left(1-\frac{1}{\operatorname{Nm}(\mathfrak{p})}\right)\geq f^{2}\prod_{p|f}\left(1-\frac{1}{p}\right)^{2}=\varphi(f)^{2},$ (6.13)
where the first product is over prime ideals of $\mathcal{O}_{\!\Delta_{0}}$ and the second is over rational primes. Since $\sqrt{\Delta}=f\Delta_{0}\equiv 0\,\left(\operatorname{mod}\,f\mathcal{O}_{\!\Delta_{0}}\right)$, the multiplicative group $\left(\mathcal{O}_{\Delta}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\cong\left(\mathbb{Z}/f\mathbb{Z}\right)^{\times}$, so $\left|\left(\mathcal{O}_{\Delta}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|=\varphi(f)$. Thus,
$\varphi(f)\leq\frac{\left|\left(\mathcal{O}_{\!\Delta_{0}}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}{\left|\left(\mathcal{O}_{\Delta}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}\leq\frac{f^{2}}{\varphi(f)}.$ (6.14)
Using (on both sides) the inequality $\varphi(f)>\frac{f}{\ell(f)}$ where $\ell(f)$ is a real-valued function satisfying $\ell(n)>1$ for all $n$ and $\ell(n)=e^{\gamma}\log\log n$ as $n\to\infty$ [34, Thm. 2.9], we have
$\frac{f}{\ell(f)}<\frac{\left|\left(\mathcal{O}_{\!\Delta_{0}}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}{\left|\left(\mathcal{O}_{\Delta}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}\leq f\ell(f),$ (6.15)
which after taking logarithms becomes
$\log f-\log\ell(f)<\log\!\left(\frac{\left|\left(\mathcal{O}_{\!\Delta_{0}}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}{\left|\left(\mathcal{O}_{\Delta}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}\right)<\log f+\log\ell(f).$ (6.16)
Using (6.7) and combining (6.11) and (6.16), for $\Delta_{0}>N$, we obtain
$\left|\log h_{\Delta}-\left(\log\sqrt{\Delta}_{0}+\log f\right)\right|<\epsilon\log\sqrt{\Delta}_{0}+\left|\log\log\varepsilon_{\Delta}\right|+\log\ell(f).$ (6.17)
On the left-hand side, $\log\sqrt{\Delta}_{0}+\log f=\log\sqrt{\Delta}\sim\log d$ as $d\to\infty$. On the right-hand side, as $d\to\infty$, $\log\sqrt{\Delta_{0}}<\log\sqrt{\Delta}\ll\log d$, while
$\log\log\varepsilon_{\Delta}\leq\log\log\varepsilon_{\Delta}^{+}=\log\log\frac{d-1+\sqrt{\Delta}}{2}\ll\log\log d,$ (6.18)
and $\log\ell(f)\ll\log\log\log d$. Thus,
$\left|\log h_{\Delta}-\log d\right|\ll\epsilon\log d\mbox{ as }d\to\infty.$ (6.19)
In the case when $\Delta_{0}\leq N$, combining (6.11) and (6.17), we instead obtain
$\left|\log h_{\Delta}-\left(\log\sqrt{\Delta}_{0}+\log f\right)\right|<B+\left|\log\log\varepsilon_{\Delta}\right|+\log\ell(f),$ (6.20)
which by a similar argument also implies
$\left|\log h_{\Delta}-\log d\right|\ll\epsilon\log d\mbox{ as }d\to\infty.$ (6.21)
Combining (6.19) and (6.21) with (6.6) proves the theorem in all cases. ∎
Corollary 6.4.
Assuming 5.4, as $d\to\infty$,
$\log\left|{\rm WHSIC}(d)/\operatorname{PEC}(d)\right|=\log d+o(\log d).$ (6.22)
Proof.
Follows from 6.3. ∎
7. Conjectures on counting Galois multiplets of Weyl–Heisenberg SICs
Recall that 4.10 implies that there are a finite number of well-defined Galois multiplets of SICs in dimension $d$, each consisting of a finite number of geometric equivalence classes; we will now refine this prediction to a precise count in every dimension. Our Multiplet Count Conjecture (7.1) predicts the number of multiplets in dimension $d$. The Order-to-Multiplet Conjecture (7.2) predicts the number of geometric equivalence classes in each multiplet.
7.1. Statement of the Multiplet Count Conjecture
[…]
Conjecture 7.1 (Multiplet Count Conjecture).
Fix a positive integer $d\neq 3$. Let $\Delta=\Delta_{d}=(d+1)(d-3)$ and $K=\mathbb{Q}(\sqrt{\Delta})$. The number of Galois multiplets of Weyl-Heisenberg SICs in dimension $d$ equals the number of quadratic orders $\mathcal{O}^{\prime}$ with $\mathcal{O}_{\!\Delta}\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{K}$. That is,
$\left|{\rm WHSIC}(d)/\approx_{GPEC}\right|=\left|\{\mathcal{O}^{\prime}:\,\mathcal{O}_{\!\Delta}\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{K}\}\right|.$ (7.1)
(Note that, if we write $\Delta=f^{2}\Delta_{0}$ for a fundamental discriminant $\Delta_{0}$, with $f>0$, then
|$\left|\{\mathcal{O}^{\prime}:\,\mathcal{O}_{\!\Delta}\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{K}\}\right|=\sigma_{0}(f),$ (7.2)
where $\sigma_{0}(f)$ denotes the number of positive divisors of $f$.)
The content of 7.1 is that the multiplets of Weyl–Heisenberg SICs are in one-to-one correspondence with orders between $\mathcal{O}$ and $\mathcal{O}_{K}$. The fact that the number of such orders is simply $\sigma_{0}(f)$ uses the fact that an order in a quadratic field is determined uniquely by its discriminant, which can be $(f^{\prime})^{2}\Delta_{0}$ for a fundamental discriminant $\Delta_{0}$ and a positive integer $f^{\prime}$. The order $\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}}$ contains $\mathcal{O}_{f^{2}\Delta_{0}}=\mathcal{O}_{\!\Delta}$ if and only if $f^{\prime}\mid f$. (These facts are not true for general number fields.)
7.2. Statement of the Extended Order-to-Multiplet Conjecture
We next formulate a conjecture positing a correspondence between orders and multiplets with certain properties. The number of geometric equivalence classes $[S]$ inside a Galois multiplet $[[S]]$ should be equal to the class number of the corresponding orders, and the fields $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$ associated to different multiplets are related in a natural way. (Recall that we have already shown these fields to be are well-defined on $\operatorname{PEC}(d)$-equivalence classes, and they are well-defined on Galois super-equivalence classes under the assumption of 4.10.)
Conjecture 7.2 (Extended Order-to-Multiplet Conjecture).
Fix a positive integer $d\neq 3$, let $\Delta=\Delta_{d}=(d+1)(d-3)$, and write $\Delta=f^{2}\Delta_{0}$ for a fundamental discrimiant $\Delta_{0}$ and some positive integer $f$. Then there is a bijection $\mathcal{M}$,
{positive divisors of $\{\mbox{positive divisors of $f$}\}\xrightarrow{\mathcal{M}}\{[[S]]:S\mbox{ a Weyl--Heisenberg line-SIC in }\mathbb{C}^{d}\}$ (7.3)
having the following properties:
  • (1)
    The number of geometric equivalence classes $[S]$ in the multiplet $\mathcal{M}(f^{\prime})=[[S]]$ is $\left|\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\right|$;
  • (2)
    If $f_{1}\mid f_{2}\mid f$, $[S_{1}]\in\mathcal{M}(f_{1})$, and $[S_{2}]\in\mathcal{M}(f_{2})$, then $F^{\rm trip}\!\left(S_{1}\right)\subseteq F^{\rm trip}\!\left(S_{2}\right)$ and $F^{\rm vec}\!\left(S_{1}\right)=F^{\rm proj}\!\left(S_{1}\right)\subseteq F^{\rm vec}\!\left(S_{2}\right)=F^{\rm proj}\!\left(S_{2}\right)$.
Part (1) of this conjecture is consistent with 5.4. The strengthening of this conjecture compared with 1.2 is that part (2) asks for additional compatibility under inclusion of the number fields $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$ with divisibility of the associated conductors $f^{\prime}$ (i.e., reverse inclusion of the associated quadratic orders). Assuming 1.1, we have $F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)$.
This conjecture can be tested by study of numerical invariants. One may attempt to reverse-engineer the map $\mathcal{M}$ from numerical data.
  1. (1)
    One first verifies that the number of multiplets matches the number of orders (7.1).
  2. (2)
    One can compute all the class numbers of the orders $\mathcal{O}$, and one can numerically compute the number of $\operatorname{PEC}$-equivalence classes in each multiplet. These multisets of numbers must be identical for a map to exist.
  3. (3)
    If all the class numbers of the orders $\mathcal{O}^{\prime}$ are pairwise distinct, then the putative map $\mathcal{M}$ is uniquely specified. Otherwise, the map exists if and only if the number of $\operatorname{PEC}(d)$-orbits in the different Galois equivalence classes are a permutation of these class numbers.
  4. (4)
    The reverse-inclusion requirement of the triple product and projection SIC-fields as the orders vary provides a further test of the conjecture.
The inclusion relations between the different fields gives further constraints on existence of such a map. One can study weaker invariants, the degrees of the fields, which are easier to compute. A necessary condition for inclusion of fields is a divisibility relation between their field degrees over $\mathbb{Q}$.
Ambiguity may remain. A case where there would be unresolvable ambiguity using numerical invariants is any case where there are two distinct multiplet $[[S_{1}]]$ and $[[S_{2}]]$ having
  1. (1)
    identical sizes $\left|[[S_{1}]]\right|=\left|[[S_{2}]]\right|$,
  2. (2)
    identical triple product fields $F^{\rm trip}\!\left(S_{1}\right)=F^{\rm trip}\!\left(S_{2}\right)$,
  3. (3)
    identical projection fields $F^{\rm proj}\!\left(S_{1}\right)=F^{\rm proj}\!\left(S_{2}\right)$.
Such an ambiguity can occur, as discovered by Markus Grassl for $d=47$ and $d=67$; see Section 9.3.
In Section 8, we present a refined 8.4 about the fields $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$, identifying them with certain ray class fields of the orders.
7.3. Verification of cases of the Multiplet Count Conjecture
We have verified that the two sets in the Multiplet Count Conjecture (7.1) have the same size for dimensions $d\leq 90$ ($d\neq 3$) under the assumption that the published lists of Weyl–Heisenberg SICs in these dimensions are complete.
Scott [38] gives a putatively complete list of numerical SICs in all dimensions $d\leq 90$, building on earlier work of Scott and Grassl [39]. Scott and Grassl find SIC fiducials by numerically minimizing a frame potential—a function whose global minima are SICs if SICs exist. Their search is constrained to Weyl–Heisenberg covariant sets of vectors. By repeating their search with a large number of seed vectors, they establish a high degree of numerical confidence in their results but do not rigorously rule out the existence of other SICs (Weyl–Heisenberg covaraint or otherwise).
7.4. Evidence for the Extended Order-to-Multiplet Conjecture
Some evidence for the Extended Order-to-Multiplet Conjecture (7.2) is given in Table 3 at the end of the paper, in dimensions $d\leq 15$. In the dimensions where there is more than one multiplet ($d\in\{7,8,11,12,15\}$), a conjectural correspondence between divisors $f^{\prime}\mid f$ and Galois multiplets (written using Scott and Grassl’s orbit labels) has been given. Under this correspondence, note that for divisors $f_{1},f_{2}$ of $f$ such that $f_{1}\mid f_{2}$, the corresponding degrees $\left.[F^{\rm proj}\!\left(S_{1}\right):K]\middle|[F^{\rm proj}\!\left(S_{2}\right):K]\right.$, and similarly for the triple product fields in the cases that have been computed. The table also supports the stronger 8.5, which implies 7.2.
8. Conjectures on precise fields of definition of Weyl–Heisenberg SICs
In this section, we give a refinement of both 5.4 and of the “class field hypothesis” of Appleby, Flammia, McConnell, and Yard [10]. The refined conjecture predicts the triple product field and the projector field associated to each multiplet.
8.1. Ray class fields of orders
The fields attached to a Weyl–Heisenberg SIC are believed to be abelian extensions of $K=\mathbb{Q}(\sqrt{\Delta_{d}})$. Class field theory gives an abstract description of abelian extensions of $K$, viewed with respect to the ideals of the maximal order $\mathcal{O}_{K}$. 7.2 implies that a Weyl–Heisenberg SIC can unambiguously be assigned an order $\mathcal{O}$ of $K$ with $\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}_{K}$. There is a long history of work which supplies a notion of class field theory attached to orders of number fields. In particular, the ring class field of an order $\mathcal{O}$ has Galois group over $K$ isomorphic to $\operatorname{Cl}(\mathcal{O})$ and generalizes the Hilbert class field, which is the maximal unramified extension of $K$ and has class group $\operatorname{Cl}(\mathcal{O}_{K})$.
In [33], the authors modify the standard constructions of class field theory to describe a distinguished collection of ray class fields of orders $H^{\mathcal{O}}_{\mathfrak{m},\Sigma}$, each specified by a level datum $(\mathcal{O};\mathfrak{m},\Sigma)$, where $\mathcal{O}$ is a (possibly non-maximal) order of $K$, $\mathfrak{m}$ is an integral $\mathcal{O}$-ideal, and $\Sigma$ is any subset of real places of $K$. In the case that $K$ is a real quadratic field, there are two infinite places, so $\Sigma$ is one of $\emptyset$, $\{\infty_{1}\}$, $\{\infty_{2}\}$, or $\{\infty_{1},\infty_{2}\}$. For the datum $(\mathcal{O};\mathcal{O},\emptyset)$, the associated field $H^{\mathcal{O}}_{\mathcal{O},\emptyset}$ is the ring class field associated to the order. For the datum $(\mathcal{O}_{K};\mathfrak{m},\Sigma)$, the associated field $H^{\mathcal{O}_{K}}_{\mathfrak{m},\Sigma}$ is the standard (Takagi) ray class field $H_{\mathfrak{m},\Sigma}$. For any fixed order $\mathcal{O}$, any abelian extension of $K$ as a subfield of some $H^{\mathcal{O}}_{\mathfrak{m},\Sigma}$.
We recall [33, Thm. 1.1]:
Theorem 8.1.
Let $K$ be a number field and $(\mathcal{O};\mathfrak{m},\Sigma)$ a level datum for $K$. There exists a unique abelian Galois extension $H_{\mathfrak{m},\Sigma}^{\mathcal{O}}/K$ with the property that a prime ideal $\mathfrak{p}$ of $\mathcal{O}_{K}$ that is coprime to the quotient ideal $\left(\mathfrak{m}:\mathcal{O}_{K}\right)$ splits completely in $H_{\mathfrak{m},\Sigma}^{\mathcal{O}}/K$ if and only if $\mathfrak{p}\cap\mathcal{O}=\pi\mathcal{O}$, a principal prime $\mathcal{O}$-ideal having $\pi\in\mathcal{O}$ with $\pi\equiv 1\,\left(\operatorname{mod}\,\mathfrak{m}\right)$ and $\rho(\pi)>0$ for $\rho\in\Sigma$.
To understand 8.1, recall that for any integral ideal $\mathfrak{m}$ of $\mathcal{O}$, the quotient ideal $\left(\mathfrak{m}:\mathcal{O}_{K}\right)$ is the largest ideal of $\mathcal{O}_{K}$ contained in $\mathfrak{m}$. An important invariant of an order $\mathcal{O}$ of an algebraic number field $K$ is its (absolute) conductor $\mathfrak{f}=\mathfrak{f}(\mathcal{O})=\left(\mathcal{O}:\mathcal{O}_{K}\right)$, which is the (set-theoretically) largest integral ideal $\mathfrak{f}\subseteq\mathcal{O}$ of $\mathcal{O}$ that is also an ideal of the maximal order $\mathcal{O}_{K}$. The conductor ideal $\mathfrak{f}(\mathcal{O})$ encodes information on all the non-invertible ideals in the order $\mathcal{O}$. The quotient ideal $\left(\mathfrak{m}:\mathcal{O}_{K}\right)$ always satisfies the inclusions (as $\mathcal{O}_{K}$-ideals)
$\mathfrak{f}(\mathcal{O})\mathfrak{m}\subseteq\left(\mathfrak{m}:\mathcal{O}_{K}\right)\subseteq\mathfrak{f}(\mathcal{O})\cap\mathfrak{m}\mathcal{O}_{K};$ (8.1)
in particular, $\left(\mathfrak{m}:\mathcal{O}_{K}\right)\subseteq\mathfrak{f}(\mathcal{O})$. The three ideals in (8.1) have the same prime divisors: A prime ideal $\mathfrak{p}$ of $\mathcal{O}_{K}$ such that $\mathfrak{p}\supseteq\mathfrak{f}(\mathcal{O})\mathfrak{m}=\mathfrak{f}(\mathcal{O})\mathfrak{m}\mathcal{O}_{K}$ satisfies either $\mathfrak{p}\supseteq\mathfrak{f}(\mathcal{O})$ or $\mathfrak{p}\supseteq\mathfrak{m}\mathcal{O}_{K}$, and thus, $\mathfrak{p}\supseteq\mathfrak{f}(\mathcal{O})\cap\mathfrak{m}\mathcal{O}_{K}$.
There is a Galois correspondence between ray class groups of an order and ray class fields of an order. The ray class group of an order $\operatorname{Cl}_{\mathfrak{m},\Sigma}(\mathcal{O})$ for level datum $(\mathcal{O};\mathfrak{m},\Sigma)$ is defined as a quotient of certain groups of invertible fractional ideals of the order $\mathcal{O}$. The ray class field $H_{\mathfrak{m},\Sigma}^{\mathcal{O}}$ of an order $\mathcal{O}$ is associated to an appropriate $\operatorname{Cl}_{\mathfrak{m},\Sigma}(\mathcal{O})$ in such a way that $\operatorname{Gal}(H_{\mathfrak{m},\Sigma}^{\mathcal{O}}/K)\simeq\operatorname{Cl}_{\mathfrak{m},\Sigma}(\mathcal{O})$ as abelian groups. Moreover, if $(\mathcal{O};\mathfrak{m},\Sigma)$ and $(\mathcal{O}^{\prime};\mathfrak{m}^{\prime},\Sigma^{\prime})$ are level data satisfying $\mathcal{O}\subseteq\mathcal{O}^{\prime}$, $\mathfrak{m}\mathcal{O}^{\prime}\subseteq\mathfrak{m}^{\prime}$, and $\Sigma\supseteq\Sigma^{\prime}$, then there is a natural quotient map $\operatorname{Cl}_{\mathfrak{m},\Sigma}(\mathcal{O})\twoheadrightarrow\operatorname{Cl}_{\mathfrak{m}^{\prime},\Sigma^{\prime}}(\mathcal{O}^{\prime})$ and a corresponding natural field inclusion $H_{\mathfrak{m}^{\prime},\Sigma^{\prime}}^{\mathcal{O}^{\prime}}\subseteq H_{\mathfrak{m},\Sigma}^{\mathcal{O}}$.
We recall [33, Thm. 1.3]:
Theorem 8.2.
For an order $\mathcal{O}$ in a number field $K$ and any level datum $(\mathcal{O};\mathfrak{m},\Sigma)$ with associated ray class field $H_{0}:=H_{\mathfrak{m},\Sigma}^{\mathcal{O}}$, there is an isomorphism $\operatorname{Art}_{\mathcal{O}}:\operatorname{Cl}_{\mathfrak{m},\Sigma}(\mathcal{O})\to\operatorname{Gal}(H_{0}/K)$, uniquely determined by its behavior on prime ideals $\mathfrak{p}$ of $\mathcal{O}$ that are coprime to $\mathfrak{f}(\mathcal{O})\cap\mathfrak{m}$, having the property that
$\operatorname{Art}_{\mathcal{O}}([\mathfrak{p}])(\alpha)\equiv\alpha^{q}\,\left(\operatorname{mod}\,\mathfrak{P}\right),$ (8.2)
where $\mathfrak{P}$ is any prime of $\mathcal{O}_{H_{0}}$ lying over $\mathfrak{p}\mathcal{O}_{K}$, and $q=p^{j}$ is the number of elements in the finite field $\mathcal{O}/\mathfrak{p}$. For any (not necessarily prime) ideal $\mathfrak{a}$ of $\mathcal{O}$ coprime to $\mathfrak{f}(\mathcal{O})\cap\mathfrak{m}$,
$\operatorname{Art}_{\mathcal{O}}([\mathfrak{a}])=\left.\operatorname{Art}([\mathfrak{a}\mathcal{O}_{K}])\right|_{H_{0}},$ (8.3)
where $\operatorname{Art}:\operatorname{Cl}_{\left(\mathfrak{m}:\mathcal{O}_{K}\right),\Sigma}(\mathcal{O}_{K})\to\operatorname{Gal}(H_{1}/K)$ is the usual Artin map in class field theory, with $H_{1}=H_{\left(\mathfrak{m}:\mathcal{O}_{K}\right),\Sigma}^{\mathcal{O}_{K}}$ being a (Takagi) ray class field for the maximal order $\mathcal{O}_{K}$, and $H_{0}\subseteq H_{1}$.
In 8.2, the set of prime ideals coprime to $\mathfrak{f}(\mathcal{O})\cap\mathfrak{m}$ includes all but finitely many of the prime ideals of $\mathcal{O}$. The Artin map is fully defined by (8.2), which specifies the map on prime ideals of $\mathcal{O}$ coprime to $\mathfrak{f}(\mathcal{O})\cap\mathfrak{m}$, because it turns out that any ray ideal class in $\operatorname{Cl}_{\mathfrak{m},\Sigma}(\mathcal{O})$ has a representative that is coprime to $\mathfrak{f}(\mathcal{O})\cap\mathfrak{m}$ (see [33, Lem. 4.12]), and ideals of $\mathcal{O}$ that are coprime to $\mathfrak{f}$ enjoy unique factorization into prime ideals (even though arbitrary nonzero ideals of $\mathcal{O}$ do not). When $\mathcal{O}=\mathcal{O}_{K}$, (8.2) specializes to the standard definition of the Artin map in the case of a Takagi ray class field.
For a real quadratic field $K$, the infinite places in $\Sigma$ relate to the notions of “wide” and “narrow” class fields. If $\mathfrak{m}=\mathcal{O}$, the (wide) ring class field is $H_{\mathcal{O},\emptyset}^{\mathcal{O}}=H_{\mathcal{O},\{\infty_{1}\}}^{\mathcal{O}}=H_{\mathcal{O},\{\infty_{2}\}}^{\mathcal{O}}$, and the narrow ring class field is $H_{\mathcal{O},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}}$. The narrow ring class field is either equal to the wide ring class field or is a degree $2$ extension thereof. More generally, there are four ray class fields associated to the ideal $\mathfrak{m}$ as well as a fifth subfield $H_{\mathfrak{m},+}^{\mathcal{O}}$ of $H_{\mathfrak{m},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}}$.
[…]
Each of the field inclusions in this diagram is either trivial or degree $2$. In the “nondegenerate” case when all have degree $2$, the Galois group $\operatorname{Gal}\!\left(H_{\mathfrak{m},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}}/H_{\mathfrak{m},\emptyset}^{\mathcal{O}}\right)\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$. The fields $H_{\mathfrak{m},\emptyset}^{\mathcal{O}}$, $H_{\mathfrak{m},+}^{\mathcal{O}}$, and $H_{\mathfrak{m},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}}$ are all Galois over $\mathbb{Q}$, whereas in the nondegenerate case $H_{\mathfrak{m},\{\infty_{1}\}}^{\mathcal{O}}$ and $H_{\mathfrak{m},\{\infty_{2}\}}^{\mathcal{O}}$ are not. The field $H_{\mathfrak{m},\emptyset}^{\mathcal{O}}$ is totally real, $H_{\mathfrak{m},+}^{\mathcal{O}}$ is (in the nondegenerate case) totally complex and CM, and $H_{\mathfrak{m},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}}$ is (in the nondegenerate case) totally complex and not CM.
8.2. Statement of the Extended Ray Class Fields of Orders Conjecture
We conjecture an identification of fields of triple product SIC-fields and projection SIC-fields with certain class fields of orders. We recall the fact that orders of a quadratic field $K$ are completely specified by their discriminants $\Delta^{\prime}=(f^{\prime})^{2}\Delta_{0}$, where $\Delta_{0}$ is the field discriminant of $K$ (a fundamental discriminant; see the beginning of Section 7.1) and $f^{\prime}$ is a positive integer.
Definition 8.3.
Fix a positive integer $d\geq 4$, write $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ for a fundamental discriminant $\Delta_{0}$, and let $f^{\prime}$ be a positive divisor of $f$. Let $\Delta^{\prime}=(f^{\prime})^{2}\Delta_{0}$ and $\mathcal{O}^{\prime}=\mathcal{O}_{\Delta^{\prime}}$. We introduce two families of class fields parametrized by $d$ and $f^{\prime}$:
$E_{d,f^{\prime}}$ $=H_{d\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}};$ (8.5)
$\tilde{E}_{d,f^{\prime}}$ $=H_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}.$ (8.6)
Here $d^{\prime}=d$ if $d$ is odd and $d^{\prime}=2d$ if $d$ is even.
Note that $E_{d,f^{\prime}}=\tilde{E}_{d,f^{\prime}}$ if $d$ is odd, and $E_{d,f^{\prime}}\subseteq\tilde{E}_{d,f^{\prime}}$ if $d$ is even. Specifically, when $d$ is even, the index $[\tilde{E}_{d,f^{\prime}}:E_{d,f^{\prime}}]=\left|\operatorname{Cl}_{2d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})/\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|=2$ by 10.2.
Conjecture 8.4 (Extended Ray Class Fields of Orders Conjecture).
Fix a positive integer $d\geq 4$, write $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ for a fundamental discriminant $\Delta_{0}$, and let $f^{\prime}\mid f$, $f^{\prime}>0$.
Let :{positive divisors of $\mathcal{M}:\{\mbox{positive divisors of $f$}\}\to\{[[S]]:S\mbox{ a Weyl--Heisenberg line-SIC in }\mathbb{C}^{d}\}$ be any function satisfying the conditions of 7.2. If $S$ is a Weyl–Heisenberg SIC with $[S]\in\mathcal{M}(f^{\prime})$, then
$F^{\rm trip}\!\left(S\right)$ $=E_{d,f^{\prime}}$ and $F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)$ $=\tilde{E}_{d,f^{\prime}}.$ (8.7)
8.4 implies that $F^{\rm proj}\!\left(S\right)=F^{\rm eproj}\!\left(S\right)$ for any Weyl–Heisenberg SIC $S$. This is because $\tilde{E}_{d,f^{\prime}}=H^{\mathcal{O}^{\prime}}_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}$ contains the ray class field of $\mathbb{Q}$ with level datum $(\mathbb{Z};d^{\prime}\mathbb{Z},\{\infty\})$, which is simply the field $\mathbb{Q}(\zeta_{d^{\prime}})=\mathbb{Q}(\xi_{d})$ of $d^{\prime}$-th roots of unity; thus, $E_{d,f^{\prime}}(\xi_{d})=\tilde{E}_{d,f^{\prime}}$.
8.4 is a natural extension of 1.1 of Appleby, Flammia, McConnell, and Yard. Under the additional (likely true) assumption that $F^{\rm vec}\!\left(S\right)$ is closed under complex conjugation (so that $F^{\rm proj}\!\left(S\right)=F^{\rm vec}\!\left(S\right)$), their conjecture becomes equivalent to the statement that $F^{\rm proj}\!\left(S\right)=\tilde{E}_{d,1}$ for any $S$ in the minimal multiplet.
8.3. Statement of the Ideal-Class-to-SIC Conjecture
We now formulate a conjecture that refines 5.4, 7.1, and 8.4 by insisting on compatibility with the Artin map. To state this conjecture, we decompose
$\operatorname{Clm}(\mathcal{O}_{\!\Delta})=\bigsqcup_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$ (8.8)
by means of 5.3.
Conjecture 8.5 (Ideal-Class-to-SIC Conjecture).
Fix a positive integer $d>3$, let $\Delta=(d+1)(d-3)$, and write $\Delta=f^{2}\Delta_{0}$ where $\Delta_{0}$ is a fundamental discriminant of a quadratic field. Then there is a bijection
$\bigsqcup_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\xrightarrow{\Omega}\operatorname{WHSIC}(d)/\operatorname{PEC}(d)$ (8.9)
under which an ideal class $\mathcal{A}\in\operatorname{Cl}(\mathcal{O}^{\prime})$ with $\mathcal{O}^{\prime}=\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}}$ maps to a geometric equivalence class $\Omega(\mathcal{A})=[S_{\mathcal{A}}]$ with associated fields $F^{\rm trip}\!\left(S_{\mathcal{A}}\right)=E_{d,f^{\prime}}$ and $F^{\rm proj}\!\left(S_{\mathcal{A}}\right)=\tilde{E}_{d,f^{\prime}}$.
Furthermore, $\Omega$ can be chosen to be compatible with the Artin map (defined in 8.2)
$Art_{\mathcal{O}^{\prime}}:\operatorname{Cl}_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}(\mathcal{O}^{\prime})\to\operatorname{Gal}(\tilde{E}_{d,f^{\prime}}/\mathbb{Q}(\Delta_{0})),$ (8.10)
in the sense that
$[S_{\bar{\mathcal{A}}\bar{\mathfrak{B}}}]=[(\operatorname{Art}_{\mathcal{O}^{\prime}}(\mathcal{A}))(S_{\bar{\mathfrak{B}}})]\mbox{ for any }\mathcal{A},\mathfrak{B}\in\operatorname{Cl}_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}(\mathcal{O}^{\prime})$ (8.11)
and their images $\bar{\mathcal{A}},\bar{\mathfrak{B}}\in\operatorname{Cl}(\mathcal{O}^{\prime})$ under the natural quotient map.
The required compatibility with the Artin map implies that there is exactly one Galois super-equivalence class of Weyl–Heisenberg SICs—equivalently, one Galois multiplet—for each order $\mathcal{O}$ between $\mathcal{O}_{\!\Delta_{0}}$ and $\mathcal{O}_{\!\Delta}$, which is 7.1.
We formulate this hypothesis because its truth would provide a unifying principle explaining the structure of all the other conjectures; we present no mechanism to construct a candidate map $\Omega$. 8.5 would define a map $\mathcal{M}$ of the type given in 7.2 by means of the following commutative diagram,
${\displaystyle\bigsqcup_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})}$${\operatorname{WHSIC}(d)/\operatorname{PEC}(d)}${positive divisors of ${\{\mbox{positive divisors of $f$}\}}$${\{[[S]]:S\mbox{ a line-SIC in }\mathbb{C}^{d}\}}$$\scriptstyle{\Omega}$$\scriptstyle{\mathcal{M}}$ (8.12)
where the downward maps send $\mathcal{A}\in\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$ to $f^{\prime}$ and $[S]$ to $[[S]]$, respectively.
8.5 is compatible with conjectures of the first author [32] and other authors [6, 11] connecting SICs to the Stark conjectures for real quadratic fields [41], particularly as formulated by Tate [43, 44], where the Stark units carry a Galois action compatible with the Artin map.
9. Theorems on degeneration and nondegeneration of inclusions of ray class fields of orders
In this section, we prove unconditional results about when the fields $E_{d,f^{\prime}}$ introduced in 8.3 can coincide, while fixing $d$ and varying $f^{\prime}$. We prove similar results about some closely related fields. These results allows us to explain some observed phenomena about SICs and their associated number fields.
The fact that two distinct multiplets $[[S_{1}]]\neq[[S_{2}]]$ can have the same associated field $F^{\rm proj}\!\left(S_{1}\right)=F^{\rm proj}\!\left(S_{2}\right)$ was discovered experimentally by Markus Grassl [27]. This observation came as a surprise, as experimental evidence in low dimensions had pointed to multiplets being determined by the associated fields.
We determine exactly when there is an equality of class fields $E_{d,f^{\prime}}=E_{d,f^{\prime\prime}}$ (equivalently, $\tilde{E}_{d,f^{\prime}}=\tilde{E}_{d,f^{\prime\prime}}$) in the case when $f^{\prime\prime}\mid f^{\prime}$. We do not rule out such an equality for $f^{\prime}$, $f^{\prime\prime}$ incomparable in the divisibility order. (Some such cases could be ruled out by considering ramification properties of primes in these fields as extensions of $K$.) Nonetheless, our results explain the examples discovered by Grassl.
We characterize all cases where an equality $E_{d,f^{\prime}}=E_{d,f^{\prime\prime}}$ happens for some distinct $f^{\prime},f^{\prime\prime}$ satisfying $f^{\prime\prime}\mid f^{\prime}\mid f$ (with $(d+1)(d-3)=f^{2}\Delta_{0}$), purely in terms of the dimension $d$. We show that the set of such $d$ has a limiting natural density of $\frac{1}{48}$.
8.4 then predicts that, in such dimensions $d$, there will be two distinct SIC multiplets that generate the same field. This assertion is experimentally testable. A calculation of this criterion for $d\leq 500$ is compared with Grassl’s experimental data for occurrence of equality of associated fields.
9.1. Preliminary results on class groups and class fields of orders
We recall results from [33] needed to relate different ray class fields to each other and determine the relative Galois group (including whether or not it is the trivial group). We first recall a result on when there is an inclusion between ray class fields of two orders $\mathcal{O}\subseteq\mathcal{O}^{\prime}$, which is [33, Thm. 6.5].
Theorem 9.1.
For two orders $\mathcal{O}\subseteq\mathcal{O}^{\prime}$ in a number field $K$ and any level datum $(\mathcal{O};\mathfrak{m},\Sigma)$ for $\mathcal{O}$, there are inclusions of ray class fields $H_{\mathfrak{m}\mathcal{O}^{\prime},\Sigma}^{\mathcal{O}^{\prime}}\subseteq H_{\mathfrak{m},\Sigma}^{\mathcal{O}}\subseteq H_{\left(\mathfrak{m}:\mathcal{O}^{\prime}\right),\Sigma}^{\mathcal{O}^{\prime}}$. (Here, the extended ideal $\mathfrak{m}\mathcal{O}^{\prime}$ is the smallest $\mathcal{O}^{\prime}$-ideal containing $\mathfrak{m}$, and the quotient ideal $\left(\mathfrak{m}:\mathcal{O}^{\prime}\right)$ is the maximal $\mathcal{O}^{\prime}$-ideal contained in $\mathfrak{m}$.)
To determine cases of equality $H_{\mathfrak{m}\mathcal{O}^{\prime},\Sigma}^{\mathcal{O}^{\prime}}=H_{\mathfrak{m},\Sigma}^{\mathcal{O}}$, we will convert the problem via the Artin isomorphism (8.2) to a calculation of isomorphisms between ray class groups in the two orders.
We next recall an exact sequence important for the calculation of ray class groups of orders and maps between them. The exact sequence relates class groups of two different orders in terms of the behavior of various subgroups of local and global unit groups of the two orders. In order to write down the exact sequence, we first need the following definition and conventions concerning the unit groups.
Definition 9.2.
For a commutative ring with unity $R$ and an ideal $I$ of $R$, define the group
$\operatorname{U}_{I}(R):=\{\alpha\in R^{\times}:\alpha\equiv 1\,\left(\operatorname{mod}\,I\right)\}=(1+I)\cap R^{\times}.$ (9.1)
If in addition $R$ has real embeddings and $\Sigma$ is a subset of the real embeddings of $R$, define
$\operatorname{U}_{I,\Sigma}(R):=\{\alpha\in R^{\times}:\alpha\equiv 1\,\left(\operatorname{mod}\,I\right)\mbox{ and }\rho(\alpha)>0\mbox{ for }\rho\in\Sigma\}.$ (9.2)
We also use the following extension of this notation: If there is an map $\phi:R_{1}\to R_{2}$ implicit in the discussion, and $I$ is an ideal of $R_{1}$, then we let $\operatorname{U}_{I}(R_{2}):=\operatorname{U}_{\phi(I)R_{2}}(R_{2})$ and $\operatorname{U}_{I,\Sigma}(R_{2}):=\operatorname{U}_{\phi(I)R_{2},\Sigma}(R_{2})$. The follow theorem is [33, Thm. 5.4].
Theorem 9.3.
Let $K$ be a number field and $\mathcal{O}\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{K}$ be orders of $K$, and let $\mathfrak{f}=\mathfrak{f}_{\mathcal{O}^{\prime}}(\mathcal{O})$. Let $\mathfrak{m}$ be an ideal of $\mathcal{O}$, $\mathfrak{m}^{\prime}$ an ideal of $\mathcal{O}^{\prime}$ such that $\mathfrak{m}\mathcal{O}^{\prime}\subseteq\mathfrak{m}^{\prime}$, and $\Sigma^{\prime}\subseteq\Sigma\subseteq\{\mbox{embeddings }K\hookrightarrow\mathbb{R}\}$. Let ${\mathfrak{d}}$ be any $\mathcal{O}^{\prime}$-ideal such that ${\mathfrak{d}}\subseteq\left(\mathfrak{m}:\mathcal{O}^{\prime}\right)$. With the $\operatorname{U}$-groups defined as in 9.2, we have the following exact sequence.
$1\to\frac{\operatorname{U}_{\mathfrak{m}^{\prime},\Sigma^{\prime}}\!\left(\mathcal{O}^{\prime}\right)}{\operatorname{U}_{\mathfrak{m},\Sigma}\!\left(\mathcal{O}\right)}\to\frac{\operatorname{U}_{\mathfrak{m}^{\prime}}\!\left(\mathcal{O}^{\prime}/{\mathfrak{d}}\right)}{\operatorname{U}_{\mathfrak{m}}\!\left(\mathcal{O}/{\mathfrak{d}}\right)}\times\{\pm 1\}^{\left|\Sigma\setminus\Sigma^{\prime}\right|}\to\operatorname{Cl}_{\mathfrak{m},\Sigma}(\mathcal{O})\to\operatorname{Cl}_{\mathfrak{m}^{\prime},\Sigma^{\prime}}(\mathcal{O}^{\prime})\to 1$ (9.3)
9.2. Nondegeneracy of certain ray class fields in a fixed order $\mathcal{O}$ varying infinite places
In this subsection, we show nondegeneracy of the inclusion of subfields of $E_{d,f^{\prime}}=H_{d\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}$ obtained by dropping one or both of the infinite places from the level datum. It follows that $E_{d,f^{\prime}}$ is totally complex and non-CM. The field $H_{d\mathcal{O}^{\prime},\{\infty_{1}\}}^{\mathcal{O}^{\prime}}$ appears in [10] (with the notation $\mathfrak{R}_{1}$) as a field generated by SIC overlaps after making appropriate choices.
Nondegeneracy will be deduced from the following lemma. This lemma is very special to the family of ray class groups we are considering, and it is crucial to all the calculations. To not interrupt the exposition, we defer the proof to Appendix A.
Lemma 9.4.
Let $d\geq 4$, $\Delta=(d+1)(d-3)$, and $\varepsilon_{d}=\frac{d-1+\sqrt{\Delta}}{2}$. Let $\Sigma$ be a subset of the real embedding of $\mathbb{Q}(\sqrt{\Delta})$. Fix an order $\mathcal{O}$ satisfying
$\mathcal{O}_{\!\Delta}:=\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})}.$ (9.4)
Then,
$\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})=\langle\varepsilon_{d}^{3}\rangle.$ (9.5)
The following result shows nondegeneracy of class groups whilst varying the infinite places. It generalizes [32, Lem. 5.3].
Theorem 9.5.
Let $d\geq 4$, $\Delta=(d+1)(d-3)$, and $\varepsilon_{d}=\frac{d-1+\sqrt{\Delta}}{2}$. For every order
$\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})},$ (9.6)
all the maps in the diagram
${\operatorname{Cl}_{d\mathcal{O},\{\infty_{1},\infty_{2}\}}(\mathcal{O})}$${\operatorname{Cl}_{d\mathcal{O},\{\infty_{1}\}}(\mathcal{O})}$${\operatorname{Cl}_{d\mathcal{O},\{\infty_{2}\}}(\mathcal{O})}$${\operatorname{Cl}_{d\mathcal{O},\emptyset}(\mathcal{O})}$ (9.7)
are surjections with kernel $\mathbb{Z}/2\mathbb{Z}$.
Proof.
We apply the exact sequence of 9.3, with $\mathcal{O}^{\prime}=\mathcal{O}$, $\mathfrak{m}=\mathfrak{m}^{\prime}=d\mathcal{O}$, $\Sigma^{\prime}\subseteq\Sigma$, and ${\mathfrak{d}}=(d\mathcal{O}:\mathcal{O})=d\mathcal{O}$. It says that there is an exact sequence
$1\to\frac{\operatorname{U}_{d\mathcal{O},\Sigma^{\prime}}\!\left(\mathcal{O}\right)}{\operatorname{U}_{d\mathcal{O},\Sigma}\!\left(\mathcal{O}\right)}\to\frac{\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/d\mathcal{O}\right)}{\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/d\mathcal{O}\right)}\times\{\pm 1\}^{\left|\Sigma\setminus\Sigma^{\prime}\right|}\to\operatorname{Cl}_{d\mathcal{O},\Sigma}(\mathcal{O})\to\operatorname{Cl}_{d\mathcal{O},\Sigma^{\prime}}(\mathcal{O})\to 1$ (9.8)
9.4 shows for any order $\mathcal{O}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})}$ containing $\varepsilon_{d}$ that $\frac{\operatorname{U}_{d\mathcal{O},\Sigma^{\prime}}\!\left(\mathcal{O}\right)}{\operatorname{U}_{d\mathcal{O},\Sigma}\!\left(\mathcal{O}\right)}=\frac{\langle\varepsilon_{d}^{3}\rangle}{\langle\varepsilon_{d}^{3}\rangle}=1$. We deduce that
$1\to\{\pm 1\}^{\left|\Sigma\setminus\Sigma^{\prime}\right|}\to\operatorname{Cl}_{d\mathcal{O},\Sigma}(\mathcal{O})\to\operatorname{Cl}_{d\mathcal{O},\Sigma^{\prime}}(\mathcal{O})\to 1$ (9.9)
It follows that the class groups $\operatorname{Cl}_{(d)}(\mathcal{O})$, $\operatorname{Cl}_{(d)\infty_{1}}(\mathcal{O})$, $\operatorname{Cl}_{(d)\infty_{2}}(\mathcal{O})$, and $\operatorname{Cl}_{(d)\infty_{1}\infty_{2}}(\mathcal{O})$ are all distinct and have the relationships given in the theorem statement. ∎
Remark 9.6.
By 8.2, we obtain for any order $\mathcal{O}$ satisfying $\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})}$ the following inclusions of ray class fields of orders,
[…]
where each edge in the diagram denotes an extension of degree 2. We also know that $\operatorname{Gal}(H_{d\mathcal{O},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}}/H_{d\mathcal{O},\emptyset}^{\mathcal{O}})\cong\{\pm 1\}^{2}\cong(\mathbb{Z}/2\mathbb{Z})^{2}$.
9.3. Cases of equality in ray class groups and ray class fields of two orders $\mathcal{O}\subsetneq\mathcal{O}^{\prime}$
In certain cases, the ray class group of an order does not change the under extension of the order, and consequently, the correpsonding ray class fields are identical. This phenomenon can be understood using the exact sequence given in 9.3: The rightmost two groups in the exact sequence are isomorphic if and only if the leftmost two groups in the sequence are isomorphic. We quantify exactly when this degeneration phenomenon occurs for the level data appearing in our conjectures about SICs.
Theorem 9.7.
Let $d\geq 4$, $\Delta=(d+1)(d-3)$, and $\varepsilon_{d}=\frac{d-1+\sqrt{\Delta}}{2}$. Write $\Delta=f^{2}\Delta_{0}$, where $\Delta_{0}$ is a fundamental discriminant. Let $\Sigma^{\prime}\subseteq\Sigma$ be subsets of $\{\infty_{1},\infty_{2}\}$. For every containment of orders
$\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})},$ (9.11)
the surjective map
$\operatorname{Cl}_{d\mathcal{O},\Sigma}(\mathcal{O})\to\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma^{\prime}}(\mathcal{O}^{\prime})$ (9.12)
induced by extension of ideals is an isomorphism if and only if $\Sigma=\Sigma^{\prime}$ and either:
  • (1)
    $\mathcal{O}=\mathcal{O}^{\prime}$, or
  • (2)
[…]
Proof.
Let $g=[\mathcal{O}^{\prime}:\mathcal{O}]$ be the relative index of the orders and let $\Delta^{\prime}=\operatorname{disc}{\mathcal{O}^{\prime}}$, so that the two orders have the following expressions as $\mathbb{Z}$-modules:
$\mathcal{O}^{\prime}$ $=\mathbb{Z}+\frac{\Delta^{\prime}+\sqrt{\Delta^{\prime}}}{2}$ and $\mathcal{O}$ $=\mathbb{Z}+g\frac{\Delta^{\prime}+\sqrt{\Delta^{\prime}}}{2}.$ (9.13)
A direct computation shows that the conductor ideal $\left(\mathcal{O}:\mathcal{O}^{\prime}\right)=g\mathcal{O}^{\prime}$, so the colon ideal $\left(d\mathcal{O}:\mathcal{O}^{\prime}\right)=dg\mathcal{O}^{\prime}$. We apply the exact sequence of 9.3, choosing $\mathfrak{m}=d\mathcal{O}$, $\mathfrak{m}^{\prime}=d\mathcal{O}^{\prime}$, $\Sigma^{\prime}\subseteq\Sigma$, and ${\mathfrak{d}}=dg\mathcal{O}^{\prime}$, to obtain
$1\to\frac{\operatorname{U}_{d\mathcal{O}^{\prime},\Sigma^{\prime}}(\mathcal{O}^{\prime})}{\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})}\to\frac{\operatorname{U}_{d\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}/dg\mathcal{O}^{\prime}\right)}{\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/dg\mathcal{O}^{\prime}\right)}\times\{\pm 1\}^{\left|\Sigma\setminus\Sigma^{\prime}\right|}\to\operatorname{Cl}_{d\mathcal{O},\Sigma}(\mathcal{O})\to\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma^{\prime}}(\mathcal{O}^{\prime})\to 1$ (9.14)
By 9.4, $\operatorname{U}_{d\mathcal{O}^{\prime},\Sigma^{\prime}}(\mathcal{O}^{\prime})=\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})=\langle\varepsilon_{d}^{3}\rangle$, so the leftmost term in the sequence is the trivial group. Thus, the rightmost map (which is the surjection induced by extension of ideals mentioned in the theorem statement) is an isomorphism if and only if the group $\frac{\operatorname{U}_{d\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}/df\mathcal{O}^{\prime}\right)}{\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/df\mathcal{O}^{\prime}\right)}\times\{\pm 1\}^{\left|\Sigma\setminus\Sigma^{\prime}\right|}=1$. We compute the order of this group as
$\left|\frac{\operatorname{U}_{d\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}/dg\mathcal{O}^{\prime}\right)}{\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/dg\mathcal{O}^{\prime}\right)}\times\{\pm 1\}^{\left|\Sigma\setminus\Sigma^{\prime}\right|}\right|=\left|{\frac{\operatorname{U}_{d\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}/dg\mathcal{O}^{\prime}\right)}{\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/dg\mathcal{O}^{\prime}\right)}}\right|\cdot\left|\{\pm 1\}^{\left|\Sigma\setminus\Sigma^{\prime}\right|}\right|=\left|{\frac{\operatorname{U}_{d\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}/dg\mathcal{O}^{\prime}\right)}{\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/dfg\mathcal{O}^{\prime}\right)}}\right|\cdot 2^{\left|\Sigma\setminus\Sigma^{\prime}\right|}.$ (9.15)
Thus, the map is an isomorphism if and only if $\operatorname{U}_{d\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}/dg\mathcal{O}^{\prime}\right)=\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/dg\mathcal{O}^{\prime}\right)$ and $\Sigma=\Sigma^{\prime}$. We now determine the cases when the former happens.
We wish to compare the group $U^{\prime}:=\operatorname{U}_{d\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}/dg\mathcal{O}^{\prime}\right)$ to its subgroup $U:=\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/dg\mathcal{O}^{\prime}\right)$ and determine when they are equal. They are clearly equal when $g=1$, giving (1); from now on, we assume $g>1$. The structure of the proof is as follows:
  • (A)
    If $g$ is an odd prime, we will show that $U^{\prime}\neq U$.
  • (B)
    If $g=2$, we will characterize exactly when $U^{\prime}=U$, giving (2).
  • (C)
    If $g=4$, we will use (B) to show that $U^{\prime}\neq U$ by considering intermediate orders between $\mathcal{O}$ and $\mathcal{O}^{\prime}$.
  • (D)
    We will use (A) and (C) to conclude that $U^{\prime}\neq U$ for all $g>2$ by considering intermediate orders between $\mathcal{O}$ and $\mathcal{O}^{\prime}$.
Let $\omega=\frac{\Delta^{\prime}+\sqrt{\Delta^{\prime}}}{2}$, so $\mathcal{O}^{\prime}=\mathbb{Z}+\omega\mathbb{Z}$ and $\mathcal{O}=\mathbb{Z}+g\omega\mathbb{Z}$. Let
$M^{\prime}$ =$=\{\beta\in\mathcal{O}^{\prime}/dg\mathcal{O}^{\prime}:\beta\equiv 1\,\left(\operatorname{mod}\,d\mathcal{O}^{\prime}\right)\}\mbox{ and}$ (9.16)
$M$ =$=\{\beta\in\mathcal{O}/dg\mathcal{O}^{\prime}:\beta\equiv 1\,\left(\operatorname{mod}\,d\mathcal{O}\right)\}.$ (9.17)
Then $U^{\prime}$ consists of the invertible elements of $M^{\prime}$, and $U$ consists of the invertible elements of $M$, that is, the elements of each that are coprime to $dg$.
The elements of $M^{\prime}$ and $M$ may be written down as follows, where $\overline{\alpha}$ stands for the equivalence class $\,\left(\operatorname{mod}\,dg\mathcal{O}^{\prime}\right)$ represented by $\alpha$.
$M^{\prime}$ =$=\left\{\overline{1+dm+dn\omega}:0\leq m,n<g\right\},\mbox{ and}$ (9.18)
$M$ =$=\left\{\overline{1+dm}:0\leq m<g\right\}.$ (9.19)
Case (A). Suppose $g=p$ is an odd prime. For $n\in\mathbb{Z}$ we have, using the norm in $\mathcal{O}_{K}$,
$\operatorname{Nm}(1+dn\omega)$ $=\left(1+dn\frac{\Delta^{\prime}+\sqrt{\Delta^{\prime}}}{2}\right)\left(1+dn\frac{\Delta^{\prime}-\sqrt{\Delta^{\prime}}}{2}\right)$ (9.20)
$=1+dn\Delta^{\prime}+d^{2}n^{2}\frac{(\Delta^{\prime})^{2}-\Delta^{\prime}}{4}.$ (9.21)
The quadratic polynomial $F(x)=1+d\Delta^{\prime}x+d^{2}\frac{(\Delta^{\prime})^{2}-\Delta^{\prime}}{4}x^{2}$ has at most two roots in $\mathbb{Z}/p\mathbb{Z}$, and $F(0)=1$ (so $x=0$ is not a root), so for $p\geq 5$ there exists some $n\in\{1,\ldots,p-1\}$ such that $\operatorname{Nm}(1+dn\omega)=F(n)\not\equiv 0\,\left(\operatorname{mod}\,p\right)$.
[…]
Finally $\operatorname{Nm}(1+dn\omega)\equiv 1\,\left(\operatorname{mod}\,d\right)$, so we conclude $\operatorname{Nm}(1+dn\omega)$ is a unit modulo $dg$, and thus $\overline{1+dn\omega}\in U^{\prime}\setminus U$, so $U\subsetneq U^{\prime}$.
Case (B). Suppose $g=2$. The elements of $M^{\prime}$ and $M$ may be written down as follows.
$M^{\prime}$ =$=\{\overline{1},\overline{1+d},\overline{1+d\omega},\overline{1+d+d\omega}\},\mbox{ and}$ (9.22)
$M$ =$=\{\overline{1},\overline{1+d}\}.$ (9.23)
Clearly, every element of $M^{\prime}$ and $M$ are coprime to $d$. If $d$ is even, then they are all coprime to $2d$, so $M^{\prime}=U^{\prime}$, $M=U$, and $U^{\prime}\neq U$. Now suppose that $d$ is odd; then $U=\{\overline{1}\}$. We compute
$\operatorname{Nm}(1+d\omega)=\begin{cases}1\,\left(\operatorname{mod}\,2d\right)&\mbox{if }\Delta^{\prime}\equiv 0\,\left(\operatorname{mod}\,8\right),\\ 1+d\,\left(\operatorname{mod}\,2d\right)&\mbox{if }\Delta^{\prime}\equiv 1\,\left(\operatorname{mod}\,8\right),\\ 1+d\,\left(\operatorname{mod}\,2d\right)&\mbox{if }\Delta^{\prime}\equiv 4\,\left(\operatorname{mod}\,8\right),\\ 1\,\left(\operatorname{mod}\,2d\right)&\mbox{if }\Delta^{\prime}\equiv 5\,\left(\operatorname{mod}\,8\right).\end{cases}$ (9.24)
and
$\operatorname{Nm}(1+d+d\omega)=\begin{cases}1+d\,\left(\operatorname{mod}\,2d\right)&\mbox{if }\Delta^{\prime}\equiv 0\,\left(\operatorname{mod}\,8\right),\\ 1+d\,\left(\operatorname{mod}\,2d\right)&\mbox{if }\Delta^{\prime}\equiv 1\,\left(\operatorname{mod}\,8\right),\\ 1\,\left(\operatorname{mod}\,2d\right)&\mbox{if }\Delta^{\prime}\equiv 4\,\left(\operatorname{mod}\,8\right),\\ 1\,\left(\operatorname{mod}\,2d\right)&\mbox{if }\Delta^{\prime}\equiv 5\,\left(\operatorname{mod}\,8\right).\end{cases}$ (9.25)
Thus,
$U^{\prime}=\begin{cases}\{\overline{1},\overline{1+d\omega}\}&\mbox{if }\Delta^{\prime}\equiv 0\,\left(\operatorname{mod}\,8\right),\\ \{\overline{1}\}&\mbox{if }\Delta^{\prime}\equiv 1\,\left(\operatorname{mod}\,8\right),\\ \{\overline{1},\overline{1+d+d\omega}\}&\mbox{if }\Delta^{\prime}\equiv 4\,\left(\operatorname{mod}\,8\right),\\ \{\overline{1},\overline{1+d\omega},\overline{1+d+d\omega}\}&\mbox{if }\Delta^{\prime}\equiv 5\,\left(\operatorname{mod}\,8\right).\end{cases}$ (9.26)
[…]
Case (C). Suppose $g=4$. Consider the intermediate order $\mathcal{O}_{1}=\mathbb{Z}+2\omega\mathbb{Z}$. We have
$[\operatorname{Cl}_{d,\Sigma}(\mathcal{O}):\operatorname{Cl}_{d,\Sigma}(\mathcal{O}^{\prime})]=[\operatorname{Cl}_{d,\Sigma}(\mathcal{O}):\operatorname{Cl}_{d,\Sigma}(\mathcal{O}_{1})][\operatorname{Cl}_{d,\Sigma}(\mathcal{O}_{1}):\operatorname{Cl}_{d,\Sigma}(\mathcal{O}^{\prime})],$ (9.27)
but $\operatorname{disc}(\mathcal{O}_{1})$ is even, and thus $[\operatorname{Cl}_{d,\Sigma}(\mathcal{O}):\operatorname{Cl}_{d,\Sigma}(\mathcal{O}_{1})]=2$ by (B). Thus, $\operatorname{Cl}_{d,\Sigma}(\mathcal{O})\neq\operatorname{Cl}_{d,\Sigma}(\mathcal{O}^{\prime})$ in this case (i.e., $U^{\prime}\neq U$).
Case (D). Now consider any $g>2$. Write $g=h_{1}h_{2}$ where either $h_{1}=4$ or $h_{1}$ is an odd prime. There is then an intermediate order $\mathcal{O}_{1}=\mathbb{Z}+h_{1}\omega\mathbb{Z}$ with $[\mathcal{O}^{\prime}:\mathcal{O}_{1}]=h_{1}$, and by (A) and (C), $[\operatorname{Cl}_{d,\Sigma}(\mathcal{O}_{1}):\operatorname{Cl}_{d,\Sigma}(\mathcal{O}^{\prime})]>1$. Thus, as in (C), $\operatorname{Cl}_{d,\Sigma}(\mathcal{O})\neq\operatorname{Cl}_{d,\Sigma}(\mathcal{O}^{\prime})$ (i.e., $U^{\prime}\neq U$). ∎
We obtain the following criterion for equality of the associated ray class fields.
Theorem 9.8.
The set of $d\geq 4$ such that, for $\Delta=(d+1)(d-3)$ and $\varepsilon_{d}=\frac{d-1+\sqrt{\Delta}}{2}$, there exist distinct orders $\mathcal{O}\subseteq\mathcal{O}^{\prime}$ with
$\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})},$ (9.28)
[…]
Proof.
[…]
The Galois correspondence gives an isomorphism
$\operatorname{Gal}\!\left(H_{d\mathcal{O},\Sigma}^{\mathcal{O}}/H_{d\mathcal{O}^{\prime},\Sigma}^{\mathcal{O}^{\prime}}\right)\cong\ker\!\left(\operatorname{Cl}_{d\mathcal{O},\Sigma}(\mathcal{O})\to\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right).$ (9.29)
Thus, $H_{d\mathcal{O},\Sigma}^{\mathcal{O}}=H_{d\mathcal{O}^{\prime},\Sigma}^{\mathcal{O}^{\prime}}$ if and only if the (always surjective) map $\operatorname{Cl}_{d\mathcal{O},\Sigma}(\mathcal{O})\to\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})$ is an isomorphism. The corollary now follows. ∎
We now summarize the degeneration phenomenon in the context and notation of the fields $E_{d,f^{\prime}}$, $\tilde{E}_{d,f^{\prime}}$ appearing in our conjectures.
Corollary 9.9.
Let $d\geq 4$. Write $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ for $\Delta_{0}$ a fundamental discriminant, and suppose $f^{\prime},f^{\prime\prime}$ are positive integers satisfying $f^{\prime\prime}\mid f^{\prime}\mid f$ and $f^{\prime}\neq f^{\prime\prime}$. The following are equivalent:
  • (1)
    rad(Δ)1(mod $\operatorname{rad}(\Delta)\equiv 1\,\left(\operatorname{mod}\,8\right)$ and $f^{\prime}=2f^{\prime\prime}$.
  • (2)
[…]
  • (3)
    $E_{d,f^{\prime}}=E_{d,f^{\prime\prime}}$
  • (4)
    $\tilde{E}_{d,f^{\prime}}=\tilde{E}_{d,f^{\prime\prime}}$
Proof.
The equivalence of (1), (2), and (3) follows immediately from 9.7 and the proof of 9.8, taking $\mathcal{O}=\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}}$ and $\mathcal{O}^{\prime}=\mathcal{O}_{(f^{\prime\prime})^{2}\Delta_{0}}$.
The equivalence of (3) and (4) follows from the fact that
$\tilde{E}_{d,f^{\prime}}^{\prime}$ $=E_{d,f^{\prime}}$ and $\tilde{E}_{d,f^{\prime\prime}}$ $=E_{d,f^{\prime\prime}}$ (9.30)
[…]
9.4. Rarity of degeneration for ray class fields of orders
We determine the density of $d$ for which the criterion of 9.7 holds.
Proposition 9.10.
We have
[…]
Proof.
[…]
Let $\mathcal{D}=\bigcup_{k\geq 0}{\mathcal{D}}_{k}$ where
$\mathcal{D}_{k}=\{d\geq 4:2^{2k}\mid\mid\Delta_{1}\mbox{ and }\Delta_{1}\equiv 1\,\left(\operatorname{mod}\,8\right)\}.$ (9.32)
We show that each $\mathcal{D}_{k}$ has a natural density
${\bf d}_{k}:=\lim_{n\to\infty}\frac{1}{N}\#\{d\in\mathcal{D}_{k}:\,d\leq N\}$ (9.33)
by showing it is a finite union of complete arithmetic progressions. The analysis below gives
${\bf d_{0}=\bf d_{1}=\bf d_{2}=\bf d_{3}=\bf d_{4}=\bf d_{5}}=0$ (9.34)
and ${\bf d_{k}}=\frac{2}{2^{2k}}$ for $k\geq 6$.
  1. (1)
[…]
  2. (2)
[…]
    1. (a)
[…]
    2. (b)
[…]
    3. (c)
[…]
    4. (d)
      The pattern continues for higher $k$, yielding a density set $\frac{1}{2^{2k+3}}$ set at level $k$.
  3. (3)
[…]
    1. (a)
[…]
    2. (b)
      The pattern continues for higher $k$, yielding a density set $\frac{1}{2^{2k+3}}$ set at level $k$.
[…]
$\sum_{k=1}^{\infty}\frac{1}{2^{5+2k}}=\frac{1}{96}.$
[…]
Remark 9.11.
The rarity of degeneration is special to the case of SIC fields. Degeneration is common in ring class towers over real quadratic fields; indeed, for any nonsquare $\Delta>0$, any prime $p$, and any fixed modulus $\mathfrak{m}$, it may be shown that the tower of fields
$H_{\mathfrak{m}}^{\mathcal{O}_{0}}\subseteq H_{\mathfrak{m}}^{\mathcal{O}_{1}}\subseteq H_{\mathfrak{m}}^{\mathcal{O}_{2}}\subseteq\cdots\subseteq H_{\mathfrak{m}}^{\mathcal{O}_{n}}\subseteq\cdots$ (9.35)
associated to the orders $\mathcal{O}_{n}=\mathbb{Z}[p^{n}\sqrt{\Delta}]$ is eventually constant. The lack of degeneration for inclusions of SIC fields seems due to the particular properties of the unit group of the base field modulo $d$. The tendency of SIC fields to be large, and “aligned SIC” towers incorporating SIC fields in different dimensions [5] to be noncyclotomic (and thus necessarily “horizontal,” in the sense of introducing infinitely many primes into both the degree and the modulus), holds for similar reasons.
9.5. Experimental testing of degeneration for ray class fields of orders
One can compare the prediction of 9.8 with experimental data on algebraic SICs, which can detect equality of SIC-fields for SICs in different Galois multiplets (which conjecturally correspond to different orders.) Only three cases of equality have been found experimentally, all by Marcus Grassl in unpublished work [27]. Those occur in dimensions
$d\in\{47,67,259\},$ (9.36)
and in each case, there is a secondary multiplet having the same projection SIC-field as the minimal multiplet (conjecturally corresponding to $\mathcal{O}^{\prime}=\mathcal{O}_{K}$ or $f^{\prime}=1$). Algebraic SICs have not been computed in every dimension up to $d=259$, but rather, Grassl has computed them is certain dimensions (including $d=259$) where the projection field has small enough expected degree to permit the calculation.
It is much easier to write down a list of dimensions $d$ and corresponding order conductors $f^{\prime},2f^{\prime}$ where our conjectures predict degeneration of SIC-fields, that is, where the conditions of 9.8 are satisfied. For $d\leq 500$, equality of ray class fields occurs for
$d\in\{47,67,83,175,211,259,303,339,431,447,467\}.$ (9.37)
This data is consistent with Grassl’s data for SIC-fields. For every such $d$, the conductor pair $(1,2)$ degenerates (i.e., gives the same field ray class field in the manner of 9.8), and an additional conductor pair degenerates in three of these dimensions. The data on ray class field degeneration for $d\leq 500$ is given in Table 2.
$d$ $\Delta_{0}$ $f^{\prime},2f^{\prime}$ $h$ $[E_{d,f^{\prime}}:\mathbb{Q}(\Delta_{0})]$
$47$ $33$ $1,2$ $1$ $1472$
$67$ $71$,2121,21 , 2 $1$ $1452$
$83$ $051$,2121,21 , 2 $2$ $9184$
$175$ $473$ $1,2$ $3$ $43200$
$211$ $689$ $1,2$ $4$ $117600$
$259$ $65$ $1,2$ $2$ $31104$
$303$ $57$ $1,2$ $1$ $40800$
 $d$ $\Delta_{0}$ $f^{\prime},2f^{\prime}$ $h$ $[E_{d,f^{\prime}}:\mathbb{Q}(\Delta_{0})]$ $303$ $75$,105105,105 , 10 $6$ $244800$ $339$ $7851$,2121,21 , 2 $8$ $408576$ $431$ $321$ $1,2$ $3$ $371520$ $431$ $213$,6363,63 , 6 $9$ $1114560$ $447$ $777$ $1,2$ $4$ $355200$ $467$ $377$ $1,2$ $2$ $290784$ $467$ $773$,6363,63 , 6 $8$ $1163136$
Table 2. SIC dims. $d\leq 500$ with an equality of ray class fields of orders $E_{d,f^{\prime}}=E_{d,f^{\prime\prime}}$. The columns show the fund. disc. $\Delta_{0}$, each pair of conductors $(f^{\prime},f^{\prime\prime})=(f^{\prime},2f^{\prime})$ exhibiting field equality, the order class number $h=h_{(f^{\prime})^{2}\Delta_{0}}=h_{(2f^{\prime})^{2}\Delta_{0}}$, and the ray class field extension degree $[E_{d,f^{\prime}}:\mathbb{Q}(\Delta_{0})]=[E_{d,2f^{\prime}}:\mathbb{Q}(\Delta_{0})]$.
10. Comparison of degrees of ray class fields of orders and degrees of SIC-fields
We give formulas for the degrees of the ray class fields of orders $E_{d,f^{\prime}}$ and $\tilde{E}_{d,f^{\prime}}$. These degrees conjecturally coincide with the SIC fields $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$ for those Weyl–Heisenberg SICs $S$ assigned to these orders in 8.4, and we make a numerical comparison in Section 11.
Definition 10.1.
If $K$ is a field and $n\in\mathbb{N}$, define the number field totient function by
$\varphi_{K}(n)=\left|\left(\mathcal{O}_{K}/n\mathcal{O}_{K}\right)^{\times}\right|.$ (10.1)
This function is a natural generalization of Euler’s totient function $\varphi(n)=\varphi_{\mathbb{Q}}(n)$. It is easily seen to be multiplicative by the Chinese remainder theorem. It may be further generalized to $\varphi_{\mathcal{O}}(\mathfrak{a}):=\left|(\mathcal{O}/\mathfrak{a})^{\times}\right|$, for any order $\mathcal{O}$ and any nonzero $\mathcal{O}$-ideal $\mathfrak{a}$, which we do not need here.
Theorem 10.2.
Let $d\geq 4$ and let $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ for some fundamental discriminant $\Delta_{0}$. Let $K=\mathbb{Q}(\sqrt{\Delta})$. Let $\varepsilon_{d}=\frac{(d-1)+\sqrt{\Delta}}{2}$, and let $\mathcal{O}^{\prime}$ be a $K$-order such that $\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{K}$, that is, with $\operatorname{disc}(\mathcal{O}^{\prime})=(f^{\prime})^{2}\Delta_{0}$ for some $f^{\prime}\mid f$. Then
$\left|\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|=\begin{cases}\dfrac{2^{\left|\Sigma\right|}h_{K}\operatorname{Reg}_{K}\varphi_{K}(df^{\prime})}{9\log(\varepsilon_{d})\varphi(f^{\prime})},&\mbox{if}d\equiv 3\,\left(\operatorname{mod}\,9\right)\mbox{ and }3\mid f^{\prime},\vspace{5pt}\\ \dfrac{2^{\left|\Sigma\right|}h_{K}\operatorname{Reg}_{K}\varphi_{K}(df^{\prime})}{6\log(\varepsilon_{d})\varphi(f^{\prime})},&\mbox{otherwise}.\end{cases}$ (10.2)
If $d$ is even, then $\left|\operatorname{Cl}_{2d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|=2\left|\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|$.
Proof.
Let $\tilde{d}=d$ if $d$ is odd and $\tilde{d}\in\{d,2d\}$ if $d$ is even. By [33, Thm. 5.6], the cardinality of the ray class group is
$\left|\operatorname{Cl}_{\tilde{d}\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|=\frac{h_{K}}{\left[\mathcal{O}_{K}^{\times}:\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right]}\cdot\frac{2^{\left|\Sigma\right|}\left|\left(\mathcal{O}_{K}\middle/\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)\right)^{\times}\right|}{\left|\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}\middle/\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)\right)\right|}.$ (10.3)
By 9.4, $\operatorname{U}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})=\langle\varepsilon_{d}^{3}\rangle$. On the other hand, when $\tilde{d}=2d$,
[…]
$\left[\mathcal{O}_{K}^{\times}:\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right]=\begin{cases}\left[\langle-1,\varepsilon_{K}\rangle:\langle\varepsilon_{d}^{3}\rangle\right]=\dfrac{2\log(\varepsilon_{d}^{3})}{\log(\varepsilon_{K})}=\dfrac{6\log(\varepsilon_{d})}{\operatorname{Reg}_{K}},&\mbox{if }\tilde{d}=d,\vspace{5pt}\\ \left[\langle-1,\varepsilon_{K}\rangle:\langle\varepsilon_{d}^{6}\rangle\right]=\dfrac{2\log(\varepsilon_{d}^{6})}{\log(\varepsilon_{K})}=\dfrac{12\log(\varepsilon_{d})}{\operatorname{Reg}_{K}},&\mbox{if }\tilde{d}=2d.\end{cases}$ (10.5)
The quotient ideal appearing in (10.3) is $\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)=\tilde{d}\left(\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)=\tilde{d}f^{\prime}\mathcal{O}_{K}$. Thus,
$\left|\left(\mathcal{O}_{K}\middle/\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)\right)^{\times}\right|=\left|\left(\mathcal{O}_{K}\middle/\tilde{d}f^{\prime}\mathcal{O}_{K}\right)^{\times}\right|=\varphi_{K}(\tilde{d}f^{\prime}).$ (10.6)
Now we must compute the size of $\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}\middle/\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)\right)=\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}\middle/\tilde{d}f^{\prime}\mathcal{O}_{K}\right)$. Write $\mathcal{O}_{K}=\mathbb{Z}+\omega\mathbb{Z}$ and $\mathcal{O}^{\prime}=\mathbb{Z}+f^{\prime}\omega\mathbb{Z}$. Thus, $\mathcal{O}^{\prime}/\tilde{d}f^{\prime}\mathcal{O}_{K}=\frac{\mathbb{Z}+f^{\prime}\omega\mathbb{Z}}{\tilde{d}f^{\prime}\mathbb{Z}+\tilde{d}f^{\prime}\omega\mathbb{Z}}$ and $\tilde{d}\mathcal{O}^{\prime}=\tilde{d}\mathbb{Z}+\tilde{d}f^{\prime}\omega\mathbb{Z}$, so the multiplicative group
$\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}\middle/\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)\right)$ =$=\left\{u\in\frac{\mathbb{Z}+f^{\prime}\omega\mathbb{Z}}{\tilde{d}f^{\prime}\mathbb{Z}+\tilde{d}f^{\prime}\omega\mathbb{Z}}:u\equiv 1\,\left(\operatorname{mod}\,\tilde{d}\mathbb{Z}+\tilde{d}f^{\prime}\omega\mathbb{Z}\right)\right\}$ (10.7)
$\cong\left\{u\in\mathbb{Z}/\tilde{d}f^{\prime}\mathbb{Z}:u\equiv 1\,\left(\operatorname{mod}\,\tilde{d}\right)\right\}.$ (10.8)
[…]
$\left|\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}\middle/\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)\right)\right|=\begin{cases}\frac{3}{2}\varphi(f^{\prime}),&\mbox{if }d\equiv 3\,\left(\operatorname{mod}\,9\right)\mbox{ and }3\mid f^{\prime},\\ \varphi(f^{\prime}),&\mbox{otherwise}.\end{cases}$ (10.9)
In the case $\tilde{d}=d$, plugging in (10.5), (10.6), and (10.9) to (10.3) yields (10.2). In the case $\tilde{d}=2d$ (with $d$ even), we have $\varphi_{K}(\tilde{d}f^{\prime})=4\varphi_{K}(df^{\prime})$, and again using plugging in (10.5), (10.6), and (10.9) to (10.3) and comparing to (10.2), we see that $\left|\operatorname{Cl}_{2d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|=2\left|\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|$. ∎
To use 10.2 for computation, the following elementary lemma is helpful.
Lemma 10.3.
Let $K$ be a quadratic field of discriminant $\Delta_{0}$. Then
$\varphi_{K}(n)=n^{2}\prod_{p|n}\left(1-\frac{1}{p}\right)\left(1-\left(\frac{\Delta_{0}}{p}\right)\frac{1}{p}\right),$ (10.10)
where the product is over primes dividing $n$, and $\left(\frac{\Delta_{0}}{p}\right)$ is the Kronecker symbol.
Proof.
The function $\varphi_{K}(n)$ is multiplicative by the Chinese remainder theorem. It is computed on prime powers by splitting into cases based on the splitting behavior of $p$ in $K$, which is measured by the Kronecker symbol. A straightforward calculation yields (10.10). ∎
Remark 10.4.
Applying the Brauer–Siegel formula to 10.2, one obtains (for $\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta_{0}})}$, $\Delta_{0}=\operatorname{disc}(\mathbb{Q}(\varepsilon_{d}))$, $\operatorname{disc}(\mathcal{O}^{\prime})=(f^{\prime})^{2}\Delta_{0}$) the asymptotic estimate
$\log\left|\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|\sim\frac{1}{2}\log((f^{\prime})^{2}\Delta_{0})+2\log(d)\mbox{ as }d\to\infty.$ (10.11)
In particular,
$d^{2+o(1)}\ll\left|\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|\ll d^{3+o(1)},$ (10.12)
with both extremes being achieved for infinitely many pairs $(d,f^{\prime})$.
11. Numerical evidence for the conjectures
We present formal numerical evidence for the conjectures, based on the published tables of Scott and Grassl [39, 38], a paper of Appleby, Chien, Flammia, and Waldron [7], and Flammia’s online tables [22]. The tables of Scott and Grassl contain putatively complete lists of $\operatorname{EC}(d)$-orbits of Weyl–Heisenberg covariant SICs as numerical approximations for $d\leq 90$ and as exact algebraic solutions for $d\leq 14$. The paper [7] produced putatively complete lists of $\operatorname{EC}(d)$-orbits as exact algebraic solutions for all $d\leq 21$. (All of these works also produced many SICs in higher dimensions, without producing a complete list.)
  • For dimensions $d\leq 90$, we check that the size of $\operatorname{Clm}(\mathbb{Z}[\varepsilon_{d}])$ is equal to the number of known $\operatorname{PEC}(d)$-orbits (or geometric equivalence classes) of Weyl–Heisenberg covariant SICs in dimension $d$, as per 5.4. This data was tabulated in Table 1 in Section 5.3.
  • For dimensions $d\leq 15$, we check that the known $\operatorname{PEC}(d)$-orbits and may be assigned to orders $\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{K}$ in such a way that $\left|\operatorname{Cl}(\mathcal{O}^{\prime})\right|$ orbits are assigned to $\mathcal{O}^{\prime}$, and those orbits form a Galois multiplet, as per 7.2. In addition, the degrees of the class fields $E_{d,f^{\prime}}$ and $\tilde{E}_{d,f^{\prime}}$ over $K=\mathbb{Q}(\sqrt{\Delta_{d}})$ match the degrees of $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$, respectively, as per 8.4. The data is tabulated in Table 3 below.
Isomorphisms of fields have not been checked except in a few small dimensions, due to computational difficulties.
Other sources of evidence also point toward the truth of these conjectures but have not been formally tabulated here. The connection to Stark units supports 8.5, including the compatibility with the Artin map [32, 6, 11]. In all cases tested (including all $d\leq 15$), the cardinality of the subgroup of $\operatorname{PEC}(d)$ fixing the line $\mathbb{C}v$ spanned by a fiducial vector $v$ (called ${\rm S}$ in [39, Table I]) equals the cardinality of the image of corresponding order unit group $(\mathcal{O}^{\prime})^{\times}$ in $(\mathcal{O}^{\prime}/d\mathcal{O}^{\prime})^{\times}$. Finally, work in preparation of Appleby, Flammia, and Kopp will present further evidence and refinements for the conjectures [8].
Orbit labels in the “multiplet” column are based on Flammia’s online table [22], following [39, 38, 7], specifically using the labels for exact SICs, which differ in some cases from Scott and Grassl’s labels for numerical SICs. We recall the definitions of the ray class fields appearing in the table:
  1. (1)
    $E_{d,f^{\prime}}:=H_{d\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}$, the ray class field having level datum $(\mathcal{O}^{\prime};d\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\})$, where $(d+1)(d-3)=f^{2}\Delta_{0}$ for a fundamental discrimiant $\Delta_{0}$, $f^{\prime}$ is a positive divisor of $f$, and $\mathcal{O}^{\prime}$ is the quadratic order of discriminant $(f^{\prime})^{2}\Delta_{0}$.
  2. (2)
    $\tilde{E}_{d,f^{\prime}}:=H_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}$, where $d^{\prime}=d$ if $d$ is odd and $d^{\prime}=2d$ if $d$ is even.
8.4 predicts that $F^{\rm trip}\!\left(S\right)=E_{d,f^{\prime}}$ and $F^{\rm proj}\!\left(S\right)=\tilde{E}_{d,f^{\prime}}$.
$d$ $\mathcal{O}^{\prime}$ $f^{\prime}$ $h_{\mathcal{O}^{\prime}}$ $[E_{d,f^{\prime}}:K]$ $[\tilde{E}_{d,f^{\prime}}:K]$ multiplet $[F^{\rm trip}\!\left(S\right):K]$ $[F^{\rm proj}\!\left(S\right):K]$
$4$ $\mathbb{Z}[\frac{1+\sqrt{5}}{2}]$ $1$ $1$ $4$ $8$ 4a $4$ $8$
$5$ $\mathbb{Z}[\sqrt{3}]$ $1$ $1$ $16$ $16$ 5a $16$ $16$
$6$ $\mathbb{Z}[\frac{1+\sqrt{21}}{2}]$ $1$ $1$ $12$ $24$ 6a $12$ $24$
$7$ $\mathbb{Z}[\sqrt{2}]$ $1$ $1$ $12$ $12$ 7b $12$ $12$
$\mathbb{Z}[2\sqrt{2}]$ $2$ $1$ $24$ $24$ 7a $24$ $24$
$8$ $\mathbb{Z}[\frac{1+\sqrt{5}}{2}]$ $1$ $1$ $8$ $16$ 8b $8$ $16$
$\mathbb{Z}[\frac{1+3\sqrt{5}}{2}]$ $3$ $1$ $32$ $64$ 8a $32$ $64$
$9$ $\mathbb{Z}[\sqrt{15}]$ $1$ $2$ $72$ $72$ 9ab $72$ $72$
$10$ $\mathbb{Z}[\frac{1+\sqrt{77}}{2}]$ $1$ $1$ $48$ $96$ 10a $48$ $96$
$11$ $\mathbb{Z}[\sqrt{6}]$ $1$ $1$ $80$ $80$ 11c $\ast$ $80$
$\mathbb{Z}[2\sqrt{6}]$ $2$ $2$ $160$ $160$ 11ab $\ast$ $160$
$12$ $\mathbb{Z}[\frac{1+\sqrt{13}}{2}]$ $1$ $1$ $16$ $32$ 12b $\ast$ $32$
$\mathbb{Z}[\frac{1+3\sqrt{13}}{2}]$ $3$ $1$ $48$ $96$ 12a $\ast$ $96$
$13$ $\mathbb{Z}[\sqrt{35}]$ $1$ $2$ $192$ $192$ 13ab $\ast$ $192$
$14$ $\mathbb{Z}[\frac{1+\sqrt{165}}{2}]$ $1$ $2$ $144$ $288$ 14ab $\ast$ $288$
$15$ $\mathbb{Z}[\sqrt{2}]$ $1$ $1$ $48$ $48$ 15d $\ast$ $48$
$\mathbb{Z}[2\sqrt{2}]$ $2$ $1$ $96$ $96$ 15b $\ast$ $96$
$\mathbb{Z}[4\sqrt{2}]$ $4$ $2$ $192$ $192$ 15ac $\ast$ $192$
Table 3. Table of $d$, orders $\mathcal{O}^{\prime}=\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}}$ with $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$, class numbers $h_{\mathcal{O}^{\prime}}$, degrees of fields $E_{d,f^{\prime}}$ and $\tilde{E}_{d,f^{\prime}}$, and index of fields $[F^{\rm trip}\!\left(S\right):K]$ and $[F^{\rm proj}\!\left(S\right):K]$. Entries denoted by $\ast$ have not been computed.
12. Concluding Remarks
This paper considered fields that are geometric invariants of a SIC. It can be useful in practice to define SICs using data in smaller subfields that are not geometric invariants. One is the field generated by the entries of a single fiducial projector $\Pi$. That field can be smaller than the full projector field by nearly a factor of $d$ is some special cases, and this is extremely useful for computing exact SICs in some large dimensions that would otherwise be inaccessible to numerical investigation [6, 11]. Another is the field generated by the SIC overlaps $\operatorname{Tr}(\Pi D_{p})$ (or else the closely related SIC overlap phases $\sqrt{d+1}\operatorname{Tr}(\Pi D_{p})$), where $D_{p}$ are the displacement operators.
The overlap field (and the overlap phase field) appear to be ray class fields with ramification at just one of the infinite places $\infty_{1}$, provided one chooses a “strongly centered” fiducial projector; see [9, 20]. That phenomenon relates to the connection to Stark units, which live in a ray class field with ramification at the other infinite place $\infty_{2}$ and are (at least in some cases) Galois conjugate to squares of overlap phases.
Work in preparation by Appleby, Flammia, and Kopp [8] will make the classification scheme suggested by this paper explicit, giving a conjectural construction of each presumptive equivalence class of Weyl–Heisenberg covariant SICs in dimension $d$. Specifically, an explicit construction of a candidate for a map $\Omega$ satisfying the conditions of 8.5 (our strongest conjecture) will be given.
13. Acknowledgements
Work of the first author was supported by NSF DMS grant #2302514. Work of the second author was supported by NSF DMS grant #1701576.
We are grateful to the developers of Magma and in particular David Kohel, whose software package reduced_forms (and whose package class_number, on which reduced_forms relies) we used to compute class numbers of real quadratic orders.
We thank Markus Grassl for finding numerical examples of degeneration of SIC-fields and sharing those examples with us. Professor Grassl’s insights prompted the first author to correct an originally incorrect statement and proof of 9.7.
We thank Marcus Appleby, Steven Flammia, and Gary McConnell for valuable conversations. We also acknowledge the contributions of Steve Donnelly, who independently connected SICs to orders in real quadratic number fields and shared his insights with Marcus Appleby, Steven Flammia, and Markus Grassl through informal discussions.
Appendix A Proof of 9.4
We will now prove 9.4, which we restate as A.1.
Lemma A.1.
Let $d\geq 4$, $\Delta=(d+1)(d-3)$, and $\varepsilon_{d}=\frac{1}{2}(d-1+\sqrt{\Delta})$. Let $\Sigma$ be a subset of the two real embeddings of $\mathbb{Q}(\sqrt{\Delta})$. Fix an order $\mathcal{O}$ satisfying
$\mathcal{O}_{\!\Delta}:=\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})}.$ (A.1)
Then,
$\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})=\langle\varepsilon_{d}^{3}\rangle.$ (A.2)
Proof.
Let $K=\mathbb{Q}(\sqrt{\Delta})$. Write $\Delta=f^{2}D$, $f\geq 1$, where $D$ is squarefree, so $K=\mathbb{Q}(\sqrt{D})$ and
$\mathcal{O}_{K}=\left\{\frac{1}{2}(a+b\sqrt{D}):(a,b)\in\mathbb{Z}^{2},\,a\equiv b\,\left(\operatorname{mod}\,2\right),\,a\equiv 0\,\left(\operatorname{mod}\,2\right)\mbox{ if }D\not\equiv 1\,\left(\operatorname{mod}\,4\right)\right\}.$ (A.3)
For the datum $(\mathfrak{m},\Sigma)$ consisting of an ideal $\mathfrak{m}$ of the order $\mathcal{O}$ in $K$ and a set of real places $\Sigma$ of $K$, recall the definition of the unit subgroup
$\operatorname{U}_{\mathfrak{m},\Sigma}(\mathcal{O}):=\left\{\alpha\in\mathcal{O}^{\times}:\alpha\equiv 1\,\left(\operatorname{mod}\,\mathfrak{m}\right),\,\sigma(\alpha)>0\mbox{ for }\sigma\in\Sigma\right\}.$ (A.4)
For A.1, the unit group of interest is $\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})$. The unit $\varepsilon_{d}$ is totally positive, and
$\varepsilon_{d}^{3}=\frac{1}{2}\left((d-1)(d^{2}-2d-2)+d(d-2)f\sqrt{D}\right)\equiv 1\,\left(\operatorname{mod}\,d\mathcal{O}_{\!\Delta}\right).$ (A.5)
Thus, we have the containment relations
$\langle\varepsilon_{d}^{3}\rangle\subseteq\operatorname{U}_{d\mathcal{O}_{\!\Delta},\{\infty_{1},\infty_{2}\}}\subseteq\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})\subseteq\operatorname{U}_{d\mathcal{O}_{K}}(\mathcal{O}_{K}).$ (A.6)
These relations show it suffices to prove (A.2) for the special case $\mathcal{O}=\mathcal{O}_{K}$ with $\Sigma=\emptyset$.
The proof studies the minimal unit $\eta>1\in\mathcal{O}_{K}$ such that
$\eta=\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right).$ (A.7)
[…]
Claim 1. We have $\operatorname{Nm}(\eta)=\eta\eta^{\prime}=1$. Consequently $\eta$ is totally positive, and $10<\eta^{\prime}<$ 1.
Proof. We know $\eta\eta^{\prime}=\pm 1$. By hypothesis, $\eta=\pm 1+d\omega$ for some $\omega\in\mathcal{O}_{K}$, and taking algebraic conjugates yields $\eta^{\prime}=\pm 1+d\omega^{\prime}$. Then $\eta\eta^{\prime}=1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$, which forces $\eta\eta^{\prime}=1$. Total positivity and the bound for $\eta^{\prime}$ follow, proving the claim.
Claim 2. For some integer $k\geq 1$, $\eta^{k}=\varepsilon_{d}^{3}.$
Proof. We have $1<\eta\leq\varepsilon_{d}^{3}$, so for some $k\geq 1$, there holds $\eta^{k-1}<\varepsilon_{d}^{3}\leq\eta^{k}$. Then $\omega:=\eta^{k}\varepsilon_{d}^{-3}$ is a unit, $1\leq\omega<\eta$, and $\omega\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$; hence by minimality of $\eta>1$, we must have $\omega=1$, proving the claim.
Claim 3. $\eta^{k}=\varepsilon_{d}^{3}$ for $k\geq 3$ cannot occur.
Proof. If $k=3$, then $\eta^{3}=\varepsilon_{d}^{3}$, which, since $\eta>1$ forces $\eta=\varepsilon_{d}$. We know $\eta\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$ and obtain a contradiction by showing that $\varepsilon_{d}\not\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$. To show the latter, we argue by contradiction and suppose $\varepsilon_{d}\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$. We know $\varepsilon_{d}\varepsilon_{d}^{\prime}=\eta\eta^{\prime}=1$, which then holds $\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$ and thus implies $\varepsilon_{d}^{\prime}\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$; hence, $\varepsilon_{d}+\varepsilon_{d}^{\prime}\equiv\pm 2\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$. However, $\varepsilon_{d}+\varepsilon_{d}^{\prime}=\frac{1}{2}(d-1+\sqrt{\Delta})+\frac{1}{2}(d-1+\sqrt{\Delta})=d-1\equiv-1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$, giving a contradiction for $d\geq 4$. Thus $\varepsilon_{d}\not\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$ holds, and $\eta\neq\varepsilon_{d}$.
If $k\geq 4$, then $\eta=(\varepsilon_{d})^{3/k}<\varepsilon_{d}$ (taking the positive $k$-th root.) Now Claim 1 gives (since $x+\frac{1}{x}$ is an increasing function for $x\geq 1$) that
$11<\eta+\eta^{\prime}=\operatorname{Tr}(\eta)<\varepsilon_{d}+\varepsilon_{d}^{\prime}=\operatorname{Tr}(\varepsilon_{d})=d-$ 1 . (A.8)
The trace is an integer, and
$\operatorname{Tr}(\eta)$ $\equiv\eta+\eta^{\prime}\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)\equiv\pm 2\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right).$ (A.9)
Since $d\geq 4$ and $\operatorname{Tr}(\eta)\in\mathbb{Z}$, the bounds on $\operatorname{Tr}(\eta)$ give
$\operatorname{Tr}(\eta)=\begin{cases}2&\mbox{if }\eta\equiv 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right),\\ d-2&\mbox{if }\eta\equiv-1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right).\end{cases}$ (A.10)
We have two cases.
  • Case 1: Suppose $\operatorname{Tr}(\eta)=2$. Then $\eta+\eta^{\prime}=2$ and $\eta\eta^{\prime}=1$, so $\eta$ is a root of the polynomial $x^{2}-2x+1=(x-1)^{2}$, so $\eta=1$, contradicting $\eta>1$.
  • Case 2: Suppose $\operatorname{Tr}(\eta)=d-2$. We may assume $d\geq 5$, since the case $d=4$ is covered by Case 1. Now
    $d-3<\eta=\operatorname{Tr}(\eta)-\eta^{\prime}<\varepsilon_{d}<\operatorname{Tr}(\varepsilon_{d})=d-1.$ (A.11)
    We have for $d\geq 6$ that
    $(d-3)^{3}<\eta^{3}<\varepsilon_{d}^{3}<(d-1)^{3}<(d-3)^{5}<\eta^{5},$ (A.12)
    since $(d-1)^{3}<(d-3)^{5}$ for $d\geq 6$. Thus $\eta^{k}=\varepsilon_{d}^{3}$ cannot hold. For $d=5$, we have $\eta+\eta^{\prime}=3$ and $\eta\eta^{\prime}=1$, so $\eta$ is a root of the polynomial $x^{3}-3x+1$, so $\eta=\frac{3+\sqrt{5}}{2}$. However, for $d=5$, $\Delta=12$, so $\varepsilon_{5}\in\mathbb{Q}(\sqrt{3})$ cannot be an integer power of $\eta$.
This completes the proof of Claim 3. It remains to rule out $k=2$, where $\varepsilon_{d}^{3}$ is the square of a unit $\tilde{\eta}$. We first classify all $d$ where such a unit exists.
Claim 4. The set of all $d\geq 4$ having a unit $\tilde{\eta}\in\mathcal{O}_{K}$ that satisfies $\tilde{\eta}^{2}=\varepsilon_{d}^{3}$ is:
  1. (1)
    $d=a^{2}-1$ for any $a\geq 3$. Then, $\tilde{\eta}=\pm\eta\varepsilon_{d}$, where
    $\eta=\varepsilon_{a+1}=\frac{a+\sqrt{a^{2}-4}}{2}$ (A.13)
    has $\operatorname{Nm}(\eta)=+1$. Here $\sqrt{\Delta}=a\sqrt{d-3}=a\sqrt{a^{2}-4}$.
  2. (2)
    $d=a^{2}+3$ for any $a\geq 1$. Then, $\tilde{\eta}=\pm\eta\varepsilon_{d}$, where $\eta=\frac{a+\sqrt{a^{2}+4}}{2}$ has $\operatorname{Nm}(\eta)=-1.$ Here $\sqrt{\Delta}=a\sqrt{d+1}=a\sqrt{a^{2}+4}$.
[…]
$\eta^{2}=\frac{1}{4}\left((a^{2}+b^{2}D)+2ab\sqrt{D}\right),$ (A.14)
while $\varepsilon_{d}=\frac{1}{2}(d-1+f\sqrt{D})$. To obtain a solution, we must have
$ab$ $=f;$ (A.15)
$a^{2}+b^{2}D$ $=2(d-1).$ (A.16)
We have $a\neq 0$ since $f\neq 0$, so the first equation gives $b^{2}=\frac{f^{2}}{a^{2}}$, and the second equation can be rewritten (since $\Delta=Df^{2}$) as
$a^{2}+\frac{\Delta}{a^{2}}=2(d-1).$ (A.17)
Introducing the new variable $y=a^{2}$, we find
$y^{2}-2(d-1)y+(d+1)(d-3)=(y-(d+1))(y-(d-3))=0,$ (A.18)
which yields two cases. The first is $y=a^{2}=d+1$, which gives solutions $d=a^{2}-1$ for $a\geq 3$, to make $d\geq 4$. The second is $y=a^{2}-3$, which gives solutions $d=a^{2}+3$ for $a\geq 1$.
  • Case 1: Suppose $d=a^{2}+1$ (so $a\geq 3$). Then $\Delta=a^{2}(a^{2}-4)$, $\sqrt{\Delta}=a\sqrt{a^{2}-4}$, and $a^{2}-4=Db^{2}$. Now
    $\eta$ $=\varepsilon_{a+1}=\frac{1}{2}(a+\sqrt{a^{2}-4});$ (A.19)
    $\varepsilon_{d}$ $=\eta^{2}=\frac{1}{2}(a^{2}-2+a\sqrt{a^{2}-4}).$ (A.20)
    Here $\operatorname{Nm}(\eta)=+1$, and the unit $\tilde{\eta}=\eta\varepsilon_{d}>1$ has $\operatorname{Nm}(\tilde{\eta})=1$.
  • Case 2: Suppose $d=a^{2}-3$, ($a\geq 1$). Then $\Delta=a^{2}(a^{2}+4)$ and $\sqrt{\Delta}=a\sqrt{a^{2}+4}$. So
    $\eta$ $=\frac{1}{2}(a+\sqrt{a^{2}+4});$ (A.21)
    $\varepsilon_{d}$ $=\eta^{2}=\frac{1}{2}(a^{2}+2+a\sqrt{a^{2}+4}).$ (A.22)
    Here $\operatorname{Nm}(\eta)=-1$, and the unit $\tilde{\eta}=\eta\varepsilon_{d}>1$ has $\operatorname{Nm}(\tilde{\eta})=-1$.
This proves Claim 4. To complete the proof, we must rule out the case $k=2$.
Claim 5. $\eta^{k}=\varepsilon_{d}^{3}$ for $k=2$ does not occur.
We must show that all the $\tilde{\eta}$ described by Claim 4 have
$\tilde{\eta}\not\equiv\pm 1\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right).$ (A.23)
The Claim 4 solutions for $d=a^{2}+1$ are ruled out by Claim 1, since $\operatorname{Nm}(\tilde{\eta})=-1$. To rule out the Claim 4 solutions having $d=a^{2}-1$, note that these solutions $\tilde{\eta}$ with $\tilde{\eta}^{2}=\varepsilon_{d}^{3}$ are given by $\tilde{\eta}=e_{a+1}e_{a^{2}-1}=e_{a+1}^{3}$ with $\varepsilon_{a+1}=\frac{1}{2}(a+\sqrt{a^{2}-4})$. We have
$\tilde{\eta}=\frac{1}{2}\left(a(a^{2}-3)+(a^{2}-1)\sqrt{a^{2}-4}\right).$ (A.24)
Now $\Delta=(d+1)(d-3)=a^{2}(a^{2}-4)$; hence, $\Delta=f^{2}D$ has $D\mid(a^{2}-4)$ and $K=\mathbb{Q}(\sqrt{a^{2}-4})$. We assert for $d=a^{2}-1$ that
$\tilde{\eta}\equiv-a\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right).$ (A.25)
The assertion implies $a\pm 1\not\in d\mathcal{O}_{K}$ holds for $d=a^{2}-1$, with $d\geq 3$, which implies Claim 5. It is proved in two cases.
[…]
    $\tilde{\eta}$ $=\frac{1}{2}(a(a^{2}-3)+(a^{2}-1)\sqrt{a^{2}-4}))$ (A.26)
    $\equiv\frac{1}{2}a(a^{2}-3)-\frac{1}{2}(a^{2}-1)$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right)$ (A.27)
    $\equiv\frac{a-1}{2}(a^{2}-1)-a$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right)$ (A.28)
    $\equiv-a$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right).$ (A.29)
  • Case 2. Suppose $a$ is even. Let $a=2a^{\prime}$. Then, $\Delta=a^{2}\Delta_{0}$ with $\Delta_{0}=a^{2}-4=4((a^{\prime})^{2}-1)$, so $K=\mathbb{Q}(\sqrt{(a^{\prime})^{2}-1})$. Now
    $\tilde{\eta}$ $=\frac{a}{2}(a^{2}-3)+(a^{2}-1)\sqrt{(a^{\prime})^{2}-1}$ (A.30)
    $\equiv\frac{a}{2}(a^{2}-1)-a$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right)$ (A.31)
    $\equiv-a$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right).$ (A.32)
We have proved the assertion, hence Claim 5, which completes the proof. ∎